### v1

# GCoT: Chain-of-Thought Prompt Learning for Graphs

Abstract.
Chain-of-thought (CoT) prompting has achieved remarkable success in natural language processing (NLP). However, its vast potential remains largely unexplored for graphs. This raises an interesting question: How can we design CoT prompting for graphs to guide graph models to learn step by step? On one hand, unlike natural languages, graphs are non-linear and characterized by complex topological structures. On the other hand, many graphs lack textual data, making it difficult to formulate language-based CoT prompting. In this work, we propose the first CoT prompt learning framework for text-free graphs, GCoT. Specifically, we decompose the adaptation process for each downstream task into a series of inference steps, with each step consisting of prompt-based inference, “thought” generation, and thought-conditioned prompt learning. While the steps mimic CoT prompting in NLP, the exact mechanism differs significantly. Specifically, at each step, an input graph, along with a prompt, is first fed into a pre-trained graph encoder for prompt-based inference. We then aggregate the hidden layers of the encoder to construct a “thought”, which captures the working state of each node in the current step. Conditioned on this thought, we learn a prompt specific to each node based on the current state. These prompts are fed into the next inference step, repeating the cycle. To evaluate and analyze the effectiveness of GCoT, we conduct comprehensive experiments on eight public datasets, which demonstrate the advantage of our approach.
Graph mining, chain-of-thought, prompt learning, pre-training, few-shot learning.
Corresponding authors.
ccs: Information systems Web miningccs: Information systems Data miningccs: Computing methodologies Learning latent representations
1. Introduction
Recently, Chain-of-Thought (CoT) prompting has demonstrated significant advancement in the field of natural language processing (NLP) (Wei et al., 2022; Wang et al., 2023a; Chu et al., 2023), which mimics the logical process a person may employ to solve a task. Instead of directly providing an answer, CoT prompting decomposes a problem into several steps, guiding the pre-trained language model to follow these steps that lead to the final answer. For example, given the math question in Fig. 1(a), a language model processes with CoT (e.g., initial quantity, eaten quantity, and subtraction) to reach the final answer of 4.
Figure 1. Illustration of chain-of-thought in NLP, standard graph prompts and graph chain-of-thought prompting.
However, the vast potential of CoT prompt learning remains unexplored on pre-trained graph models. Graphs can capture the interactions between entities in a wide range of domains, exhibiting a non-linear topological structure, such as molecular graphs (Wang et al., 2023b; Lee et al., 2023), citation networks (Kanakia et al., 2019; Xiong et al., 2017), and social networks (Ji et al., 2023; Zhang et al., 2023). Conventional approaches typically retrain graph neural networks (GNNs) (Kipf and Welling, 2017; Veličković et al., 2018) or graph transformers (Ying et al., 2021; Yun et al., 2019) for each specific task in an end-to-end supervised manner, relying on abundant labeled data. More recent pre-training methods (Velickovic et al., 2019; You et al., 2020) learn task-invariant, general properties from unlabeled graphs through self-supervised pretext tasks and are then fine-tuned via task-specific labels to adapt to various downstream applications (Hu et al., 2020a; Qiu et al., 2020). To further narrow the gap between pre-training and downstream tasks, prompt learning (Yu et al., 2024a; Liu et al., 2023; Fang et al., 2024) has emerged as a low-resource alternative. They unify the objectives of pre-training and downstream tasks using the same template and employ a lightweight prompt to modify either the input features or hidden embeddings of the pre-trained graph encoder, while keeping the pre-trained weights frozen, as shown in Fig. 1(b). A contemporary work (Jin et al., 2024) leverages graphs to guide reasoning; however, it is built on the CoT mechanism from NLP and relies on textual data to construct thoughts, precluding its application to general text-free graphs. For text-free graphs, existing graph learning methods—including supervised, pre-training, and prompting approaches—produce a “final answer” in a single inference step, which may limit the refinement of the final prediction. In this work, we explore the following question: Would introducing additional inference steps in a CoT style enhance the ability of pre-trained graph models to refine their predictions?
[…]
First, what should be the inference steps and thoughts for a graph task? In NLP, a thought is defined as a short instruction that reflects a single reasoning step, with each intermediate textual answer serving as a thought (Chu et al., 2023). However, in general text-free graphs, we cannot directly leverage text as prompts or thoughts. In this work, we aim to improve the inference capability of a pre-trained graph model by incorporating additional steps to refine the answer. We design an inference step with three substages: prompt-based inference, thought construction, and prompt learning. For prompt-based inference, we feed the input graph for the downstream task, along with the prompts, into a pre-trained graph encoder. Then, we construct a “thought” by fusing embeddings from each layer of the pre-trained encoder to capture the current working state with varying levels of graph topological knowledge (Kipf and Welling, 2017). Lastly, the thought is used to learn a set of prompts that guide the next step.
Second, how can we leverage a “thought” to learn prompts and guide the next-step inference? In NLP, chain-of-thought (CoT) prompting is typically implemented by appending specific phrases such as “let’s think step by step” or by providing few-shot CoT examples (Wei et al., 2022; Feng et al., 2024). Then, following a given prompt template, the language model generates new thoughts based on the query and prior thoughts, which in turn facilitate the next reasoning step. In our work, the absence of textual data prevents us from explicitly guiding the next step. Moreover, since each node in a graph exhibits unique characteristics, inference may benefit from node-specific prompts. Thus, we propose a thought-conditioned prompt learning method to guide the next inference step. Specifically, inspired by conditional prompt learning (Zhou et al., 2022), we generate a unique prompt for each node via a conditional network (condition-net), which is conditioned on the node-specific element of the previously constructed thought. The generated prompts then feed into the graph encoder to initiate and guide the next step, repeating the process.
In summary, the contributions of this work are fourfold. (1) We propose GCoT, a Graph CoT prompting approach to guide pre-trained graph models to perform step-by-step inference. To the best of our knowledge, this is the first exploration of CoT prompting on text-free graphs. (2) We design an inference step with three substages: prompt-based inference, thought construction, and thought-conditioned prompt learning. In particular, a thought is constructed by fusing embeddings from each layer of the graph encoder to capture fine-grained topological knowledge in each step. (3) We employ a condition-net to generate node-specific prompts based on the previous thought, enabling the model to perform inference for the downstream task through a step-by-step, node-specific adaptation. (4) We conduct extensive experiments on eight benchmark datasets, demonstrating the superior performance of GCoT compared to a suite of state-of-the-art methods.
2. Related Work
Chain-of-Thought prompting. Chain-of-Thought (CoT) prompting has emerged as a groundbreaking technique in natural language processing (NLP), empowering language models to address complex reasoning tasks by producing intermediate reasoning steps (Wei et al., 2022; Wang et al., 2023a; Gao et al., 2023). By breaking down problems into a sequence of logical steps, CoT emulates human-like thought processes, leading to significant improvements in model performance on tasks requiring structured reasoning (Yao et al., 2024; Feng et al., 2024). Despite its remarkable success in the NLP domain, the potential of CoT for graphs remains largely unexplored. A contemporary work, GraphCoT (Jin et al., 2024), leverages the inherent relational information in text-attributed graphs (Yan et al., 2023; Wen and Fang, 2023) to guide the reasoning process. However, GraphCoT primarily focuses on question-answering tasks for natural language and cannot be extended to general graphs lacking textual descriptions.
Graph representation learning. GNNs (Kipf and Welling, 2017; Veličković et al., 2018) are dominant technique for graph representations learning. They generally update nodes embeddings iteratively by aggregating information from their local neighborhoods based on a message-passing mechanism (Xu et al., 2019; Hamilton et al., 2017). Despite their success, GNNs often demand substantial amounts of task-specific labeled data and necessitate retraining for each new task, limiting their flexibility and scalability. Recently, researchers have extensively investigated pre-training techniques for graphs (Kipf and Welling, 2016; Hu et al., 2020b, a; Lu et al., 2021). These approaches involve pre-training a graph encoder using self-supervised objectives, and then adapt the pre-trained knowledge to downstream tasks. However, a significant gap exists between the objectives of pre-training and those of downstream tasks, resulting in suboptimal performance.
Graph prompt learning. First proposed in NLP, prompt learning has emerged as a powerful framework for bridging the gap between pre-training and downstream tasks (Brown et al., 2020; Liu et al., 2021b; Lester et al., 2021). Recently, this paradigm has been extended to the graph domain as a compelling alternative to fine-tuning approaches (Liu et al., 2023; Sun et al., 2022; Yu et al., 2024c, d). These methods typically utilize a universal template to align pre-training and downstream tasks, followed by task-specific prompts that facilitate seamless adaptation to downstream tasks while keeping the pre-trained model frozen. However, current graph prompt learning methods directly produce a final answer in a single step, resulting in insufficient refinement to the answer.
3. Preliminaries
Graph. A graph is defined as $G=(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges. The nodes are associated with a feature matrix $\mathbf{X}\in\mathbb{R}^{|V|\times d}$, where $\mathbf{x}_{v}\in\mathbb{R}^{d}$ is a row of $\mathbf{X}$ representing the feature vector for node $v\in V$. For a collection of multiple graphs, we define it as $\mathcal{G}=\{G_{1},G_{2},\dots,G_{N}\}$.
Figure 2. Overall framework of GCoT.
Graph encoder. Towards graph representation learning, one of the most widely used families of graph encoders is graph neural networks (GNNs), which generally rely on message passing to capture structural knowledge (Wu et al., 2020; Zhou et al., 2020). Each node updates its representation by aggregating information from its neighbors, and stacking multiple GNN layers enables iterative message propagation across the graph. Formally, let $\mathbf{H}^{l}$ denote the embedding matrix at the $l$-th layer, where each row $\mathbf{h}_{i}^{l}$ represents the embedding of node $v_{i}$. This matrix is iteratively computed using the embeddings from the preceding layer:
(1) $\mathbf{H}^{l}=\textsc{MP}(\mathbf{H}^{l-1},G;\theta^{l}),$
where $\textsc{MP}(\cdot)$ is the message passing function, and $\theta^{l}$ represents the learnable parameters of the graph encoder at layer $l$. The initial embedding matrix, $\mathbf{H}^{0}$, is the input feature matrix, i.e., $\mathbf{H}^{0}=\mathbf{X}$. The output after a total of $L$ layers is then $\mathbf{H}^{L}$; for brevity we simply write $\mathbf{H}$. We abstract the multi-layer encoding process as
(2) $\{\mathbf{H}^{1},\mathbf{H}^{2},\cdots,\mathbf{H}^{L}\}=\textsc{GraphEncoder}(\mathbf{X},G;\Theta),$
where $\{\mathbf{H}^{1},\mathbf{H}^{2},\cdots,\mathbf{H}^{L}\}$ denotes the embedding matrix of the each layer of the graph encoder, respectively. $\Theta=(\theta^{1},\ldots,\theta^{L})$ is the collection of weights across the layers.
Pre-training. As stated by prior studies (Yu et al., 2024d, 2025), all contrastive pre-training task on graphs (Liu et al., 2023; Velickovic et al., 2019; You et al., 2020) can be unified under the task template of similarity calculation. Formally, the unified pre-training objective is defined as follows:
(3) $\mathcal{L}(\Theta)=-\sum_{o\in\mathcal{T}_{\text{pre}}}\ln\frac{\sum_{a\in Pos_{o}}\exp(\text{sim}(\mathbf{h}_{a},\mathbf{h}_{o})/\tau)}{\sum_{b\in Neg_{o}}\exp(\text{sim}(\mathbf{h}_{b},\mathbf{h}_{o})/\tau)},$
where $Pos_{o}$ and $Neg_{o}$ denote the sets of positive and negative samples for a target instance $o$, respectively. $\mathbf{h}_{o}$ represents the embedding of the target instance, while $\mathbf{h}_{a}$ and $\mathbf{h}_{b}$ correspond to the embeddings of positive and negative samples. The hyperparameter $\tau$ controls the temperature scaling in the similarity computation. In our framework, we follow previous work (Liu et al., 2023; Yu et al., 2024d) by employing similarity calculation as the task template and using link prediction as the pre-training task.
Problem definition. In this work, we explore chain-of-thought (CoT) prompt learning framework for text-free graphs. We focus on two widely used tasks in graph learning: node classification and graph classification, in few-shot scenarios. For node classification, given a graph $G=(V,E)$ with a set of node classes $Y$, each node $v_{i}\in V$ is associated with a label $y_{i}\in Y$. In contrast, graph classification considers a collection of graphs $\mathcal{G}$, where each graph $G_{i}\in\mathcal{G}$ is assigned a class label $Y_{i}\in Y$. In the few-shot setting, only $m$ labeled examples per class are available (e.g., $m\leq 10$), a paradigm defined as $m$-shot classification (Liu et al., 2023; Yu et al., 2024d).
4. Chain-of-Thought Graph Prompt Learning
In this section, we propose our model, GCoT, starting with an overview of its framework. Then we detail its core components and conclude with a complexity analysis of the algorithm.
4.1. Overall framework
We illustrate the overall framework of GCoT in Fig. 2, consisting of two key stages: pre-training and CoT prompting. First, we pre-train a graph encoder as shown in Fig. 2(a). Details of pre-training are provided in Sect. 3. Second, given a pre-trained graph encoder, to guide the model take additional inference step before finalizing predictions, we propose CoT prompting, as shown in Fig. 2(c). Specifically, we design an inference step with three substages: prompt-based inference, thought construction, and thought-conditioned prompt learning. We first feed the prompt modified query graph into the pre-trained graph encoder. Then we construct a thought by fusing embeddings of all hidden layers of the pre-trained graph encoder. Conditioned on the thought from the previous step, we employ a condition-net to generate a series of node-specific conditional prompts that capture individualized learning patterns for nodes in a fine-grained and parameter-efficient manner, as shown in in Fig. 2(b). We repeat the inference steps until obtain a final answer, as shown in Fig. 2(c). Moreover, to enhance alignment between downstream tasks and the pre-training objective, we incorporate an initial prompt to adjust features or embeddings, same as prior graph prompting methods (Liu et al., 2023; Fang et al., 2024).
4.2. Chain-of-Thought Prompting
To bridge the gap between the objectives between pre-training and downstream tasks, we adhere to previous graph prompting methods, leveraging an initial prompt to modify node features (Sun et al., 2022; Fang et al., 2024) or output embeddings (Liu et al., 2023; Yu et al., 2025), as shown in Fig. 2(c). Note that our framework is fully compatible with any standard graph prompting techniques and can further enhance their performance, as demonstrated in Sect. 5.5. In our experiments, we employ an attention mechanism to generate prompts that modify the output embedding of the final step, as detailed in the Prompt tuning section.
Prompt-based inference For the $k$-th inference step, we first feed the prompt modified query graph into the pre-trained encoder:
(4) $\{\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}\}=\textsc{GraphEncoder}(\mathbf{X}_{k},G;\Theta_{0})$
where $\Theta_{0}$ is the pre-trained parameters in the graph encoder, $\mathbf{X}_{k}$ is the modified feature by prompts, which will be introduced in the Thought conditioned prompt learning section.
Thought construction To leverage the hierarchical knowledge across multiple layers of the graph encoder, we construct the thought by fusing the hidden embeddings from each layer of the pre-trained graph encoder as follows:
(5) $\mathbf{T}_{k}=\mathtt{Fuse}(\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}),$
where $\mathbf{H}_{k}^{l}$ denotes the hidden embedding from the $l$-th layer during the $k$-th inference step. The $\mathtt{Fuse}(\cdot)$ function can be implemented in various type. In our experiments, we employ weighted summation as the fusion function:
(6) $\mathbf{T}_{k}=w^{1}\cdot\mathbf{H}_{k}^{1}+w^{2}\cdot\mathbf{H}_{k}^{2}+\cdots+w^{L}\cdot\mathbf{H}_{k}^{L},$
where $w^{1},w^{2},\cdots,w^{L}$ are learnable parameters. The thought $\mathbf{T}_{k}$ is then used to guide the the next inference step.
Thought conditioned prompt learning The thought generated in the previous inference step stores hierarchical structural knowledge, which guides the learning in the subsequent step. Since different nodes may exhibit distinct characteristics with respect to the downstream task, rather than learn all node’s representation in the same manner, the pre-trained model should be steered to learn node-specific patterns. To this end, we propose leveraging a condition-net (Zhou et al., 2022; Yu et al., 2024b, 2025) to generate node-specific prompts, guiding node-specific adaptation. Specifically, conditioned on the thought from the previous step, $\mathbf{T}_{k}$, the condition-net generates a series of thought prompts as follows:
(7) $\mathbf{P}_{k}=\mathtt{CondNet}(\mathbf{T}_{k};\phi),$
where $\mathtt{CondNet}$ is the condition-net parameterized by $\phi$. Note that condition-net is a lightweight hypernetwork (Ha et al., 2022), where our condition-net $\mathtt{CondNet}$ functions as an auxiliary network to generate the prompts conditioned on the thoughts parameter-efficiently. In our implementation, we utilize a simple multi-layer perceptron (MLP) with a bottleneck structure for improved efficiency (Wu and Lee, 2018). The $i$-th row of $\mathbf{P}_{k}$ represents a unique node-specific prompt vector $\mathbf{p}_{i,k}$ for node $v_{i}$, which is derived from the thought $\mathbf{T}_{k}$. This prompt vector then guides the next inference step by modifying the node features as follows:
(8) $\mathbf{X}_{k+1}=\mathbf{P}_{k}\odot\mathbf{X},$
where $\odot$ denotes element-wise multiplication, and $\mathbf{X}_{k+1}$ represents the input to the pre-trained graph encoder in the $k+1$ step. After repeating $K$ inference steps, the final step outputs the nodes embeddings $\mathbf{H}_{K}$.
Prompt tuning For initial prompts, all standard graph prompting methods can be applied within our framework. In our experiments, we adhere to GPF+ (Fang et al., 2024) to generate initial prompts. Specifically, we train $N$ bias prompts $\{\mathbf{p}_{\text{bias}}^{1},\mathbf{p}_{\text{bias}}^{2},\ldots,\mathbf{p}_{\text{bias}}^{N}\}$, and leverage attention-based aggregation to generate node-specific prompts. However, unlike GPF+, which uses these prompts to modify graph features, we use them to modify the final embeddings. Specifically, the task prompt for node $i$ is computed as:
(9) $\mathbf{p}_{\text{init}}^{i}=\sum_{j=1}^{N}\alpha_{i,j}\mathbf{p}_{\text{bias}}^{j},\quad\text{where}\quad\alpha_{i,j}=\frac{\exp\big{(}\mathbf{a}^{j}\mathbf{h}_{K}^{i}\big{)}}{\sum_{l=1}^{N}\exp\big{(}\mathbf{a}^{l}\mathbf{h}_{K}^{i}\big{)}}.$
Here, $\mathbf{h}_{K}^{i}$ denotes the $i$-th row of the final embedding matrix $\mathbf{H}_{K}$, which serves as the embedding for node $i$, and $\mathbf{A}=\{\mathbf{a}^{1},\mathbf{a}^{2},\ldots,\mathbf{a}^{N}\}$ are $N$ learnable linear projection vectors. The generated initial prompts form the matrix $\mathbf{P}_{\text{init}}$, which is then used to modify the final output embeddings as follows:
(10) $\mathbf{\tilde{H}}=\mathbf{P}_{\text{init}}\odot\mathbf{H}_{K}.$
Note that for clarity and to maintain a unified representation of the various task prompt mechanisms, we use $\mathbf{P}_{\text{init}}$ to denote the trainable parameters associated with these mechanisms throughout the remainder of this paper.
Finally, for a given task with a labeled training set
$\mathcal{D}_{t}=\{(x_{1},y_{1}),(x_{2},y_{2}),\dots\},$
where each $x_{i}$ represents either a node or a graph and $y_{i}\in Y$ denotes its corresponding class label, the downstream loss function is defined as:
(11) $\mathcal{L}_{\text{down}}(\phi,\mathbf{P}_{\text{init}})=-\sum_{(x_{i},y_{i})\in\mathcal{D}_{t}}\ln\frac{\exp\left(\text{sim}\left(\mathbf{\tilde{h}}_{x_{i}},\mathbf{\bar{h}}_{y_{i}}\right)/\tau\right)}{\sum_{c\in Y}\exp\left(\text{sim}\left(\mathbf{\tilde{h}}_{x_{i}},\mathbf{\bar{h}}_{c}/\right)\tau\right)},$
where $\mathbf{\tilde{h}}_{x_{i}}$ represents the final embedding of a node $v$ or a graph $G$. Specifically, for node classification, $\mathbf{\tilde{h}}_{v}$ corresponds to a row in the answer matrix $\mathbf{\tilde{H}}$, while for graph classification, we compute the graph embedding as
$\mathbf{\tilde{h}}_{G}=\sum_{v\in V}\mathbf{\tilde{h}}_{v},$
incorporating an additional graph readout step. The prototype embedding for each class $c$, denoted as $\mathbf{\bar{h}}_{c}$, is obtained by averaging the embeddings of all labeled nodes or graphs belonging to that class. During prompt tuning, only the task prompts and lightweight parameters of the condition network ($\phi$) are updated, while the pre-trained GNN weights remain frozen. This parameter-efficient design makes our approach well-suited for few-shot learning, where the training set $\mathcal{D}_{t}$ contains only a limited number of labeled examples.
4.3. Algorithm and Complexity Analysis
Algorithm. We detail the main steps for Chain-of-Thought graph prompting in Algorithm 1. In lines 3–12, we iterate through $K$ inference steps during the downstream adaptation phase while keeping the pre-trained weights $\Theta_{0}$ frozen. Specifically, in lines 4–7, we generate the thought for the $k$-th inference step by fusing the output embeddings from each layer of the pre-trained graph encoder. In lines 8–12, we leverage this thought to generate node-specific prompts that guide the subsequent inference step to capture node-specific patterns. In lines 13–14, we employ a standard task prompt to modify the output embedding of the last inference step. Note that, based on the task prompt mechanism described in the original paper, these prompts can also be applied to modify the input features at the first inference step. Finally, in lines 15–17, we update the embeddings for the prototypical nodes/graphs based on the few-shot labeled data.
Algorithm 1 Chain-of-Thought Graph Prompt Learning
1:Pre-trained graph encoder with parameters $\Theta_{0}$.
2:Optimized parameters $\phi$ of condition-net, and task prompt $\mathbf{P}_{\text{init}}$.
3:while not converged do
4:    $\phi_{i}\leftarrow$ initialization
5:    $\mathbf{X}_{1}\leftarrow\mathbf{X}$
6:    /* 1 to $K-1$ inference steps*/
7:    while inference step $1\leq k<K$ do
8:         /* Encoding by the pre-trained graph encoder */
9:         $\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}\leftarrow\textsc{GraphEncoder}(G,\mathbf{X}_{k};\Theta_{0})$
10:         /* Thought construction */
11:         $\mathbf{T}_{k}\leftarrow\mathtt{Fuse}(\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L})$
12:         /* Thought prompting */
13:         /* Generate thought prompts by Eq. 7 */
14:         $\mathbf{P}_{k}\leftarrow\mathtt{CondNet}(\mathbf{T}_{k};\phi)$
15:         /* Thought prompts modification by Eq. 8 */
16:         $\mathbf{X}_{k+1}\leftarrow\mathbf{P}_{k}\odot\mathbf{X}$
17:    /* $K$-th inference step */
18:    $\{\mathbf{H}_{K}^{1},\mathbf{H}_{K}^{2},\cdots,\mathbf{H}_{K}^{L}\}\leftarrow\textsc{GraphEncoder}(G,\mathbf{X}_{K};\Theta_{0})$
19:    $\mathbf{H}_{K}\leftarrow\mathbf{H}_{K}^{L}$
20:    /* Initial prompt modification by Eq. 10*/
21:    $\mathbf{\tilde{H}}\leftarrow\mathbf{P}_{\text{init}}\odot\mathbf{H}_{K}$
22:    /* Update prototypical instance */
23:    for each class $c$ do
24:         ${\mathbf{\bar{h}}}_{c}\leftarrow\textsc{Average}(\mathbf{\tilde{h}}_{x}$: instance $x$ belongs to class $c$)
25:    /* Optimizing the parameters in condition-net */
26:    Calculate $\mathcal{L}_{\text{down}}(\phi,\mathbf{P}_{\text{init}})$ by Eq. (11)
27:    Update $\phi$ by backpropagating $\mathcal{L}_{\text{down}}(\phi,\mathbf{P}_{\text{init}})$
28:return $\{\phi,\mathbf{P}_{\text{task}}\}$
Complexity analysis. For a downstream graph $G$, we perform $K$ iterative inference steps. Each step comprises two key components: (1) thought generation using a pre-trained GNN (executed for $K$ iterations) and (2) thought-based conditional prompt learning (executed for $K-1$ iterations). The computational complexity of the first component is largely determined by the GNN architecture. In a standard GNN, each node aggregates information from up to $D$ neighboring nodes per layer. Consequently, computing node embeddings over $L$ layers incurs a complexity of $O(D^{L}\cdot|V|)$, where $|V|$ denotes the number of nodes. Aggregating the outputs from all $L$ layers adds an extra complexity of $O(L\cdot|V|)$. Since the thought generation process is repeated $K$ times, its overall complexity is:$O(K\cdot(D^{L}+L\cdot|V|))$. The second component, thought-based conditional prompt learning, refines the learning process by leveraging the thought representation from the previous step. This component consists of two stages: prompt generation and prompt tuning. In the prompt generation stage, the thought is fed into the condition-net, with a complexity of $O(|V|)$. In the prompt tuning stage, each node is adjusted via a prompt vector, also incurring a complexity of $O(|V|)$. Since this process is executed for $K-1$ iterations, the total complexity for this phase is: $O(2(K-1)\cdot|V|)$. Additionally, we employ a standard task prompt to modify node features or embeddings. The complexity of this process depends on the specific prompting mechanism. Here, we assign it a complexity of $O(|V|)$, which is common among graph prompting methods (Liu et al., 2023; Fang et al., 2024). In summary, the overall computational complexity of GCoT is: $O((D^{L}+L+2(K-1)+1)\cdot|V|)$.
4.4. Why GCoT Works
In GCoT, we adapt the following mechanism. (1) All three approaches leverage a universal task template to unify pre-training and downstream tasks, ensuring that GCoT can efficiently adapt to different downstream tasks, particularly in few-shot settings. (2) The chain-of-thought mechanism enables GCoT to perform multiple inference steps, thereby refining the final answer more effectively. (3) The thoughts in GCoT fuse hierarchical topological knowledge from graphs, allowing it to better capture structural information. (4) Conditioned on the thought from the previous step, GCoT generates a series of node-specific prompts that efficiently capture fine-grained node characteristics, thus effectively refining answer step by step. The combination of above mechanisms ensures the effectiveness of GCoT.
5. Experiments
In this section, we conduct experiments to evaluate GCoT, and analyze the empirical results.
Table 1. Summary of datasets.
Datasets Graphs Graph classes Avg. nodes Avg. edges Node features Node classes Task (N/G)
Cora 1 - 2,708 5,429 1,433 7 N
Citeseer 1 - 3,327 4,732 3,703 6 N
Pubmed 1 - 19717 88648 500 3 N
Photo 1 - 7650 238162 745 8 N
MUTAG 188 2 17.9 18.9 7 - G
COX2 467 2 41.2 43.5 3 - G
BZR 405 2 35.8 38.4 3 - G
PROTEINS 1,113 2 39.1 72.8 4 3 G
indicates the downstream task category for each dataset: ”N” corresponds to node classification, while ”G” signifies graph classification.
Table 2. Accuracy (%) evaluation on node and graph classification.
Methods Node classification Graph classification
Cora Citeseer Pubmed Photo MUTAG COX2 BZR PROTEINS
GCN 32.50±14.21 26.36±9.03 52.18±8.70 60.18±12.04 43.44±15.14 50.95±23.48 47.25±16.59 40.28±0.03
GAT 31.00±16.22 27.71±8.74 50.02±8.88 51.79±12.85 37.33±10.81 50.58±26.16 46.55±16.57 40.39±0.04
DGI/InfoGraph 54.11±9.60 45.00±9.19 57.98±11.36 64.48±10.04 53.17±17.29 53.82±14.19 49.33±15.11 52.51±10.29
GraphCL 51.96±9.43 43.12±9.61 61.18±11.60 63.93±10.21 54.92±17.09 53.81±14.21 49.73±14.66 53.81±8.97
ProG 51.35±13.92 35.08±8.63 64.07±14.74 68.85±9.99 51.99±4.50 53.45±15.01 53.52±11.97 52.73±6.57
GPF 57.35±13.87 43.04±8.85 57.61±11.26 65.26±9.89 56.55±13.95 54.16±14.07 48.65±13.96 53.05±7.62
GPF+ 64.91±15.49 42.28±8.71 57.85±10.86 69.80±10.84 56.81±12.93 55.24±13.29 50.83±19.74 54.58±8.70
GraphPrompt 54.56±12.56 46.23±8.70 64.85±10.72 68.99±9.28 55.44±12.56 54.34±14.77 54.59±10.52 53.80±7.93
GCoT 66.09±15.17 47.77±10.93 70.57±17.10 77.51±11.76 58.75±15.42 56.26±15.52 58.03±23.44 56.24±8.60
Best results are bolded and runner-up results are underlined.
5.1. Experimental Setup
Datasets. We conduct experiments on eight widely used benchmark datasets, spanning citation networks, e-commerce, protein structures, and molecular graphs. Cora (McCallum et al., 2000), Citeseer (Sen et al., 2008), and Pubmed (Sen et al., 2008) are citation networks, each consisting of a single graph. In these datasets, nodes represent academic papers and edges denote citation relationships between them. Photo (Shchur et al., 2018) is an e-commerce co-purchase network derived from Amazon’s photography-related product category, where nodes correspond to products and edges indicate frequently co-purchased items. PROTEINS (Borgwardt et al., 2005) is a dataset of protein structures. In each graph, nodes correspond to secondary structures, and edges capture spatial or sequential relationships within the amino acid sequence. MUTAG (Rossi and Ahmed, 2015), BZR (Rossi and Ahmed, 2015), and COX2 (Rossi and Ahmed, 2015) are molecular graph datasets, representing nitroaromatic compounds, ligands associated with benzodiazepine receptors, and COX2 contains molecular structures related to cyclooxygenase-2 inhibitors, respectively. A detailed summary of these datasets is presented in Table 1, with further descriptions in Appendix A.
Baselines. We compare GCoT with state-of-the-art methods across three categories: (1) Supervised GNNs: GCN (Kipf and Welling, 2017) and GAT (Veličković et al., 2018) are trained directly on downstream labels in a fully supervised manner, without any pre-training. (2) Graph pre-training models: DGI/InfoGraph [^1] DGI is originally designed for node-level tasks, while InfoGraph extends it to graph-level classification. In our experiments, we apply DGI to node classification and InfoGraph to graph classification. (Veličković et al., 2018; Sun et al., 2020) and GraphCL (You et al., 2020) adopt a “pre-train, fine-tune” strategy. These models first perform self-supervised pre-training using unlabeled graphs and are later fine-tuned for downstream tasks, where a classifier is trained with few-shot labels while keeping the pre-trained encoder frozen in single step. (3) Graph prompt learning models: ProG (Sun et al., 2023), GPF (Fang et al., 2024), GPF+ (Fang et al., 2024), and GraphPrompt (Liu et al., 2023) employ self-supervised pre-training followed by prompt tuning. Unlike the fine-tuning methods, these methods leverage a unified task template, and train task-specific prompts in single step for downstream adaptation. Further descriptions of these baselines and implementation details are shown in Appendix B &  C, respectively.
Downstream tasks and evaluation. We perform experiments on two downstream tasks: node classification and graph classification. Both tasks follow an $m$-shot classification setup, where for each class, we randomly select $m$ instances (nodes or graphs) as labeled examples. In our main results, we set $m=1$ for both node and graph classification tasks. Additionally, to examine the robustness of our method, we vary $m$ within the range $1\leq m\leq 10$, allowing us to analyze performance under different few-shot scenarios. We construct 100 independent $m$-shot tasks for each classification setting through repeated sampling. Each task is evaluated using five different random seeds, resulting in a total of 500 experimental runs per task type. We report both the mean and standard deviation across these results. We evaluate performance using accuracy, in line with previous studies (Wang et al., 2020; Liu et al., 2021a, 2023).
5.2. Performance Evaluation
We first evaluate one-shot classification tasks. Then, we examine the effect of increasing the number of shots on model performance.
One-shot performance. We present the results for one-shot node and graph classification tasks in Tables 2. We observe that: (1) GCoT consistently outperforms most baseline methods, surpassing the best competitor by margins of up to 11.05% in node classification and 6.03% in graph classification. These results underscore the advantage of incorporating multiple reasoning steps to better adapt to diverse tasks. (2) Standard graph prompt learning approaches—ProG, GPF, GPF+, and GraphPrompt—exhibit significantly weaker performance compared to GCoT. Their limitations arise from relying on a single-step inference process that lacks iterative refinement of the final answer. This contrast highlights the effectiveness of our chain-of-thought prompting, which enables step-by-step inference and captures individualized learning patterns for each node.
Few-shot performance. To evaluate the impact of labeled data availability on GCoT’s performance, we vary the number of shots in both node and graph classification tasks. The results, shown in Fig. 3, compare GCoT against several competitive baselines. First, GCoT consistently outperforms the baselines, particularly in low-shot scenarios (e.g., $m\leq 5$), where labeled data is scarce. Second, as the number of labeled samples increases (e.g., $m>5$), all methods generally exhibit improved performance, which aligns with expectations. Nevertheless, GCoT remains highly competitive, often achieving the best results. Note that on certain datasets, such as PROTEINS, performance across methods tends to show high variance. A possible reason is that this dataset exhibits greater variability in graph sizes compared to other datasets: The standard deviation of graph sizes in PROTEINS is 45.78, whereas other datasets fall within the range of 4.04 to 15.29. This may contribute to unstable performance. Despite this, GCoT demonstrates greater robustness than the competing methods. For the rest of the experiments, we focus on the one-shot classification setting.
Figure 3. Impacts of shots on node and graph classification.
5.3. Ablation Study and Visualization
[…]
Table 3. Ablation study on the effects of key components.
Methods Node classification Graph classification
Cora Pubmed MUTAG PROTEINS
[…]
GCoT-L1 61.45±15.82 59.94±14.84 56.54±14.12 54.71±8.57
GCoT-L2 66.95±16.07 63.05±15.13 57.68±13.84 54.77±8.81
GCoT-L3 64.39±15.76 69.27±18.20 57.85±16.10 56.22±8.45
GCoT 66.09±15.17 70.57±17.10 58.75±15.42 56.24±8.60
Figure 4. Visualization of output embedding space of nodes. Different colors represents different classes.
5.4. Heterophily Sensitivity
To examine the robustness of GCoT on heterophilic graphs (Pei et al., 2020; Yu et al., 2025), we conduct one-shot node classification on heterophilic datasets (Wisconsin (Pei et al., 2020) and Squirrel (Rozemberczki et al., 2021)). We provide the detailed description of these datasets in Appendix D. As shown in Table 4, GCoT consistently outperforms other competing baselines, further validating its effectiveness. These results indicates that the iterative inference process in GCoT successfully generalizes across both homophilic and heterophilic graphs.
Table 4. Accuracy (%) evaluation on heterophilic datasets.
Methods Wisconsin Squirrel
DGI 28.04 $\pm$ 6.47 20.00 $\pm$ 1.86
GraphCL 29.85 $\pm$ 8.46 21.42 $\pm$ 2.22
ProG 32.95 $\pm$ 11.42 20.61 $\pm$ 3.01
GPF 30.29 $\pm$ 10.04 19.89 $\pm$ 3.19
GPF+ 29.26 $\pm$ 9.61 20.93 $\pm$ 3.98
GraphPrompt 29.96 $\pm$ 9.40 21.22 $\pm$ 1.80
GCoT 33.04 $\pm$ 9.71 22.03 $\pm$ 4.94
5.5. Flexibility of Graph Prompting Methods
To evaluate the flexibility and robustness of GCoT, we evaluate its performance using various standard graph prompting methods as the task prompt. Specifically, we integrate ProG (Sun et al., 2023), GPF (Fang et al., 2024), GPF+ (Fang et al., 2024), and GraphPrompt (Liu et al., 2023) into our framework. The results for both node and graph classification on four datasets are presented in Table 5. Across nearly all cases, GCoT consistently outperforms its single-step prompting counterparts, regardless of the task prompt employed. These findings highlight the robustness of our approach and further confirm the advantage of chain-of-thought prompting over previous single-step prompting.
Table 5. Evaluation of GCoT with different standard prompting methods.
Prompting Thinking Node classification Graph classification
Cora Pubmed MUTAG PROTEINS
ProG - 51.35±13.92 64.07±14.74 51.99±14.50 52.73±6.57
GCoT 59.62±14.20 72.19±14.92 58.42±15.39 55.82±8.79
GPF - 57.35±13.87 57.61±11.26 56.55±13.95 53.05±7.62
GCoT 60.28±15.69 66.24±12.61 57.66±13.11 56.98±8.30
GPF+ - 64.91±15.49 57.85±10.86 56.81±12.93 54.58±8.70
GCoT 62.71±15.18 65.97±11.09 58.54±13.29 55.45±9.20
GraphPrompt - 54.56±12.56 64.85±10.72 55.44±12.56 53.80±7.93
GCoT 55.83±13.69 72.73±16.18 58.28±15.47 55.32±8.43
“-” refers to standard graph prompting without our CoT design.
5.6. Hyperparameter Analysis for Condition-Net
In our experiments, we implement the condition-net as a two-layer MLP with a bottleneck design. To examine its effect on performance, we vary the hidden dimension $s$ and present the results in Fig. 5. Our findings reveal that $s=32$ generally yields competitive results for node classification, and $s=8$ for graph classification, leading us to adopt these settings in our experiments. A smaller $s$ may restrict the model’s representational capacity, limiting its effectiveness, whereas a larger $s$ introduces additional trainable parameters and increases the likelihood of overfitting in few-shot learning scenarios.
Figure 5. Impact of hidden dimension $s$ in the condition-net.
5.7. Effect of Varying Number of Inference Steps
We further vary the number $K$ of inference steps to analyze their impact, with the results presented in Fig. 6. We observe that for both node and graph classification tasks, $K=2$ generally yields optimal performance. Hence, we adopt $K=2$ in our experiments. For Proteins, $K=3$ achieves the best result, likely due to its inherent complexity as described in Sect. 5.2.
Figure 6. Impact of number of inference steps $K$.
6. Conclusions
In this paper, we propose GCoT, the first CoT prompting framework for graphs. We define an inference step with three substages: prompt-based inference, thought construction, and thought conditioned prompt learning. Specifically, we first feed the prompt modified query into the pre-trained encoder, and then construct a thought by fusing the hidden embeddings from each layer of the pre-trained graph encoder. To guide the subsequent inference step, we generate a series of prompts conditioned on the thought from the previous step. By repeating the above inference steps, GCoT obtain the answer. Finally, we conduct extensive experiments on eight public datasets, demonstrating that GCoT significantly outperforms a range of state-of-the-art baselines.
Appendices
A. Further Descriptions of Datasets
In this section, we provide more comprehensive descriptions of the benchmark datasets used in our experiments, as summarized in Table 1, in the following.
  • Cora [^2] https://github.com/shchur/gnn-benchmark/raw/master/data/npz/cora.npz (McCallum et al., 2000) is a citation network composed of 2,708 research papers in the field of computing, each classified into one of seven categories. The network consists of 5,429 citation links between papers. Each paper is represented by a binary word vector, where each entry indicates the presence or absence of a word from a predefined vocabulary of 1,433 unique terms.
  • Citeseer [^3] https://github.com/shchur/gnn-benchmark/raw/master/data/npz/citeseer.npz (Sen et al., 2008) includes 3,312 computer science publications, categorized into six distinct classes, separate from those in Cora. The citation network contains 4,732 edges. Each document is encoded as a binary word vector that captures the presence or absence of words from a dictionary comprising 3,703 unique terms.
  • PubMed [^4] https://github.com/shchur/gnn-benchmark/raw/master/data/npz/pubmed.npz (Sen et al., 2008) is a citation network of 19,717 biomedical articles related to diabetes, divided into three categories. The network includes 44,338 citation edges. Unlike Cora and Citeseer, each document is represented by a TF/IDF-weighted word vector derived from a dictionary of 500 unique terms.
  • Photo [^5] https://github.com/shchur/gnn-benchmark/raw/master/data/npz/amazon_electronics_photo.npz (Shchur et al., 2018) consists of 7,487 photography-related products, each assigned to one of eight categories. The co-purchase network contains 119,043 edges, where connections indicate products frequently bought together. Each product is described by a feature vector extracted from metadata and customer reviews, with category labels corresponding to product types.
  • MUTAG [^6] https://huggingface.co/datasets/graphs-datasets/MUTAG (Debnath et al., 1991) is a dataset of nitroaromatic compounds aimed at predicting their mutagenic effects on Salmonella typhimurium. Each compound is modeled as a graph, where nodes represent atoms with categorical labels (encoded as one-hot vectors) based on atom types, and edges depict the chemical bonds connecting them. The dataset comprises 188 molecular graphs with 7 unique node types.
  • BZR [^7] https://chrsmrrs.github.io/datasets/docs/datasets/ (Rossi and Ahmed, 2015) consists of 405 molecular graphs representing ligands that interact with benzodiazepine receptors. Each molecule is treated as an independent graph and is classified into one of two categories.
  • COX2888https://chrsmrrs.github.io/datasets/docs/datasets/ (Rossi and Ahmed, 2015) includes 467 molecular structures of cyclooxygenase-2 inhibitors. In this dataset, nodes correspond to atoms, while edges define chemical bonds—which may be single, double, triple, or aromatic. The molecules are divided into two distinct classes.
  • PROTEINS [^9] https://chrsmrrs.github.io/datasets/docs/datasets/ (Borgwardt et al., 2005) is a dataset of protein structure graphs that encode both biochemical and structural properties. In this dataset, nodes represent secondary structural elements, while edges capture connectivity based on spatial proximity or amino acid sequence adjacency. Each node falls into one of three categories, and graphs are classified into two broader groups.
B. Further Descriptions of Baselines
In this section, we provide additional details about the baselines used in our experiments.
(1) Supervised GNNs:
  • GCN (Kipf and Welling, 2017): A graph neural network that aggregates node information using mean-pooling, thereby enabling nodes to capture structural information from their neighbors.
  • GAT (Veličković et al., 2018): Unlike GCN, GAT incorporates attention mechanisms to assign different weights to neighboring nodes, refining the aggregation process based on their relative importance.
(2) Graph Pre-training and Fine-tuning Models:
  • DGI (Veličković et al., 2018): A self-supervised pre-training method that maximizes mutual information between local node embeddings and the graph’s global representation, thereby enhancing structural awareness.
  • InfoGraph (Sun et al., 2020): An extension of DGI designed for graph-level tasks, aligning node and graph representations by optimizing their similarity.
  • GraphCL (You et al., 2020): A contrastive learning framework that leverages diverse graph augmentations to extract structural patterns, aiming to improve representation consistency across transformations.
(3) Single-step Prompting Models:
  • ProG (Sun et al., 2023): Reformulates node- and edge-level tasks as graph-level problems by employing prompt graphs with task-specific structures to guide adaptation.
  • GPF (Fang et al., 2024): A universal prompt-tuning strategy for pre-trained graph models that transforms input graph features to mimic various prompting effects.
  • GPF+ (Fang et al., 2024): An enhanced version of GPF that integrates an attention mechanism to dynamically refine prompt representations.
  • GraphPrompt (Liu et al., 2023): Bridges pre-training and downstream tasks using subgraph similarity-based prompting, where a learnable prompt is optimized to incorporate task-relevant information for both node and graph classification.
C. Implementation Details
Environment. Optimizer: The environment in which we ran all experiments is listed below.
  • Operating system: Ubuntu 22.04.2,
  • CPU information: AMD EPYC 7742 64-Core Processor,
  • GPU information: NVIDIA GeForce RTX 3090 (24 GB).
We outline key settings for the baselines and GCoT.
Baseline settings. We utilized the official codes for all open-source baselines. Each model was tuned based on the settings recommended in their respective work to achieve optimal performance.
For the baseline GCN (Kipf and Welling, 2017), we employ a 2-layer architecture, and set the hidden dimensions to 64. For GAT (Veličković et al., 2018), we employ a 2-layer architecture and set the hidden dimension to 64. Additionally, we apply 8 attention heads in the first GAT layer.
For DGI (Veličković et al., 2018), we utilize a 1-layer GCN as the backbone and set the hidden dimensions to 256. Additionally, we employ prelu as the activation function. For InfoGraph (Sun et al., 2020), a 1-layer GCN is used as the backbone, with its hidden dimensions set to 256. For GraphCL (You et al., 2020), a 1-layer GCN is also employed as its backbone, with the hidden dimensions set to 256. Specifically, we select edge dropping as the augmentations, with a default augmentation ratio of 0.2. For GraphPrompt (Liu et al., 2023), a 3-layer GCN is used as the backbone for all datasets, with the hidden dimensions set to 256. For GPF (Fang et al., 2024), we employ a 3-layer GCN as the backbone for all datasets, following the recommended settings. The hidden dimensions are set to 256. For GPF+ (Fang et al., 2024), we employ a 3-layer GCN as the backbone for all datasets, following the recommended settings. The hidden dimensions are set to 256. For ProG (Sun et al., 2023),
GCoT settings. For our proposed GCoT, we utilize a 3-layer GCN as the backbone for all datasets, with the hidden dimensions set to 256. We set the number of inference step as 2, and the number of bias prompts to generate inital prompts as 5.
D. Details about Heterophilic Datasets
To examine the robustness of GCoT on heterophilic graphs, we conduct experiments on heterophilic datasets in Sect. 5.4. The details of these datasets are as follows: (1) Cornell (Pei et al., 2020) is a network of 183 nodes representing webpages, with 295 edges indicating hyperlink connections among them. Node features are constructed using a bag-of-words approach based on webpage content. Each node is assigned to one of five categories: student, project, course, staff, or faculty. (2) Squirrel (Rozemberczki et al., 2021) consists of 5,201 Wikipedia pages discussing predefined topics. Nodes represent individual webpages, while edges capture 217,073 hyperlink connections between them, forming a page-page network. The dataset is categorized into five groups based on average monthly traffic. Node features are constructed using key informative nouns extracted from the text content of the Wikipedia pages.

### v2

@@
 To evaluate and analyze the effectiveness of GCoT, we conduct comprehensive experiments on eight public datasets, which demonstrate the advantage of our approach.
-Graph mining, chain-of-thought, prompt learning, pre-training, few-shot learning.
-Corresponding authors.
-ccs: Information systems Web miningccs: Information systems Data miningccs: Computing methodologies Learning latent representations
+Graph learning, chain-of-thought, prompt learning.
+journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1245-6/25/08doi: 10.1145/3711896.3736974ccs: Information systems Data mining
+KDD Availability Link:
+The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15501903.
 1. Introduction
-Recently, Chain-of-Thought (CoT) prompting has demonstrated significant advancement in the field of natural language processing (NLP) (Wei et al., 2022; Wang et al., 2023a; Chu et al., 2023), which mimics the logical process a person may employ to solve a task.
+Recently, Chain-of-Thought (CoT) prompting has demonstrated significant advancement in the field of natural language processing (NLP) (Wei et al., 2022; Wang et al., 2023a; Chu et al., 2023), which mimics the logical process a human may employ to solve a task.
 Instead of directly providing an answer, CoT prompting decomposes a problem into several steps, guiding the pre-trained language model to follow these steps that lead to the final answer.
 For example, given the math question in Fig.
-1(a), a language model processes with CoT (e.g., initial quantity, eaten quantity, and subtraction) to reach the final answer of 4.
+1(a), a language model utilizes CoT prompting (e.g., initial quantity, eaten quantity, and subtraction) to reach the final answer of 4.
 Figure 1. Illustration of chain-of-thought in NLP, standard graph prompts and graph chain-of-thought prompting.
 However, the vast potential of CoT prompt learning remains unexplored on pre-trained graph models.
-Graphs can capture the interactions between entities in a wide range of domains, exhibiting a non-linear topological structure, such as molecular graphs (Wang et al., 2023b; Lee et al., 2023), citation networks (Kanakia et al., 2019; Xiong et al., 2017), and social networks (Ji et al., 2023; Zhang et al., 2023).
+Graphs can capture the interactions between entities in a wide range of domains, exhibiting a non-linear topological structure, such as molecular graphs (Wang et al., 2023b; Lee et al., 2023; Li et al., 2024a), citation networks (Kanakia et al., 2019; Xiong et al., 2017; Li et al., 2024b), and social networks (Ji et al., 2023; Zhang et al., 2023).
 Conventional approaches typically retrain graph neural networks (GNNs) (Kipf and Welling, 2017; Veličković et al., 2018) or graph transformers (Ying et al., 2021; Yun et al., 2019) for each specific task in an end-to-end supervised manner, relying on abundant labeled data.
 More recent pre-training methods (Velickovic et al., 2019; You et al., 2020) learn task-invariant, general properties from unlabeled graphs through self-supervised pretext tasks and are then fine-tuned via task-specific labels to adapt to various downstream applications (Hu et al., 2020a; Qiu et al., 2020).
-To further narrow the gap between pre-training and downstream tasks, prompt learning (Yu et al., 2024a; Liu et al., 2023; Fang et al., 2024) has emerged as a low-resource alternative.
+To further narrow the gap between pre-training and downstream tasks, prompt learning (Yu et al., 2024b; Liu et al., 2023; Fang et al., 2024) has emerged as a low-resource alternative.
 They unify the objectives of pre-training and downstream tasks using the same template and employ a lightweight prompt to modify either the input features or hidden embeddings of the pre-trained graph encoder, while keeping the pre-trained weights frozen, as shown in Fig.
@@
 We design an inference step with three substages: prompt-based inference, thought construction, and prompt learning.
-For prompt-based inference, we feed the input graph for the downstream task, along with the prompts, into a pre-trained graph encoder.
-Then, we construct a “thought” by fusing embeddings from each layer of the pre-trained encoder to capture the current working state with varying levels of graph topological knowledge (Kipf and Welling, 2017).
+For prompt-based inference, we feed the input graph for the downstream task, along with some prompts, into a pre-trained graph encoder.
+Then, we construct a “thought” by fusing embeddings from each layer of the pre-trained encoder to capture the current working state with varying levels of topological knowledge (Kipf and Welling, 2017).
 Lastly, the thought is used to learn a set of prompts that guide the next step.
 Second, how can we leverage a “thought” to learn prompts and guide the next-step inference?
-In NLP, chain-of-thought (CoT) prompting is typically implemented by appending specific phrases such as “let’s think step by step” or by providing few-shot CoT examples (Wei et al., 2022; Feng et al., 2024).
+In NLP, CoT prompting is typically implemented by appending specific phrases such as “let’s think step by step” or by providing few-shot CoT examples (Wei et al., 2022; Feng et al., 2024).
 Then, following a given prompt template, the language model generates new thoughts based on the query and prior thoughts, which in turn facilitate the next reasoning step.
@@
 (1) We propose GCoT, a Graph CoT prompting approach to guide pre-trained graph models to perform step-by-step inference.
-To the best of our knowledge, this is the first exploration of CoT prompting on text-free graphs.
+To the best of our knowledge, this is the first exploration of CoT-style prompting on text-free graphs.
 (2) We design an inference step with three substages: prompt-based inference, thought construction, and thought-conditioned prompt learning.
@@
 2. Related Work
+In this section, we briefly review related work on CoT prompting, graph learning, and graph prompt learning.
 Chain-of-Thought prompting.
-Chain-of-Thought (CoT) prompting has emerged as a groundbreaking technique in natural language processing (NLP), empowering language models to address complex reasoning tasks by producing intermediate reasoning steps (Wei et al., 2022; Wang et al., 2023a; Gao et al., 2023).
+Chain-of-Thought (CoT) prompting has emerged as a groundbreaking technique in NLP, empowering language models to address complex reasoning tasks by producing intermediate reasoning steps (Wei et al., 2022; Wang et al., 2023a; Gao et al., 2023).
 By breaking down problems into a sequence of logical steps, CoT emulates human-like thought processes, leading to significant improvements in model performance on tasks requiring structured reasoning (Yao et al., 2024; Feng et al., 2024).
-Despite its remarkable success in the NLP domain, the potential of CoT for graphs remains largely unexplored.
+Despite its remarkable success in NLP applications, the potential of CoT prompting for graphs remains largely unexplored.
 A contemporary work, GraphCoT (Jin et al., 2024), leverages the inherent relational information in text-attributed graphs (Yan et al., 2023; Wen and Fang, 2023) to guide the reasoning process.
-However, GraphCoT primarily focuses on question-answering tasks for natural language and cannot be extended to general graphs lacking textual descriptions.
+However, GraphCoT primarily focuses on question-answering tasks for natural language and cannot be extended to general text-free graphs that lack textual descriptions.
 Graph representation learning.
-GNNs (Kipf and Welling, 2017; Veličković et al., 2018) are dominant technique for graph representations learning.
-They generally update nodes embeddings iteratively by aggregating information from their local neighborhoods based on a message-passing mechanism (Xu et al., 2019; Hamilton et al., 2017).
+GNNs (Kipf and Welling, 2017; Veličković et al., 2018; Yu et al., 2023; Jiang et al., 2024, 2023b) are the dominant technique for graph representations learning.
+They generally update node embeddings iteratively by aggregating information from their local neighborhoods based on a message-passing mechanism (Xu et al., 2019; Hamilton et al., 2017; Li et al., 2024c).
 Despite their success, GNNs often demand substantial amounts of task-specific labeled data and necessitate retraining for each new task, limiting their flexibility and scalability.
-Recently, researchers have extensively investigated pre-training techniques for graphs (Kipf and Welling, 2016; Hu et al., 2020b, a; Lu et al., 2021).
+Recently, researchers have extensively investigated pre-training techniques for graphs (Kipf and Welling, 2016; Hu et al., 2020b, a; Lu et al., 2021; Liu et al., 2025; Jiang et al., 2023a).
 These approaches involve pre-training a graph encoder using self-supervised objectives, and then adapt the pre-trained knowledge to downstream tasks.
@@
 First proposed in NLP, prompt learning has emerged as a powerful framework for bridging the gap between pre-training and downstream tasks (Brown et al., 2020; Liu et al., 2021b; Lester et al., 2021).
-Recently, this paradigm has been extended to the graph domain as a compelling alternative to fine-tuning approaches (Liu et al., 2023; Sun et al., 2022; Yu et al., 2024c, d).
+Recently, this paradigm has been extended to the graph domain as a compelling alternative to fine-tuning approaches (Liu et al., 2023; Sun et al., 2022; Yu et al., 2024a, c, d).
 These methods typically utilize a universal template to align pre-training and downstream tasks, followed by task-specific prompts that facilitate seamless adaptation to downstream tasks while keeping the pre-trained model frozen.
-However, current graph prompt learning methods directly produce a final answer in a single step, resulting in insufficient refinement to the answer.
+However, current graph prompt learning methods directly produce a final answer in a single step, resulting in insufficient refinement to the answer (Yu et al., 2024e, 2025a; Fang et al., 2025).
 3. Preliminaries
+In this section, we present the background and preliminaries relevant to our work.
 Graph.
@@
 The nodes are associated with a feature matrix $\mathbf{X}\in\mathbb{R}^{|V|\times d}$, where $\mathbf{x}_{v}\in\mathbb{R}^{d}$ is a row of $\mathbf{X}$ representing the feature vector for node $v\in V$.
-For a collection of multiple graphs, we define it as $\mathcal{G}=\{G_{1},G_{2},\dots,G_{N}\}$.
+For a collection of multiple graphs, we denote it as $\mathcal{G}=\{G_{1},G_{2},\dots,G_{N}\}$.
 Figure 2. Overall framework of GCoT.
@@
 The initial embedding matrix, $\mathbf{H}^{0}$, is the input feature matrix, i.e., $\mathbf{H}^{0}=\mathbf{X}$.
-The output after a total of $L$ layers is then $\mathbf{H}^{L}$; for brevity we simply write $\mathbf{H}$.
+The output after a total of $L$ layers is then $\mathbf{H}^{L}$; for brevity we may simply write $\mathbf{H}$.
 We abstract the multi-layer encoding process as
 (2) $\{\mathbf{H}^{1},\mathbf{H}^{2},\cdots,\mathbf{H}^{L}\}=\textsc{GraphEncoder}(\mathbf{X},G;\Theta),$
-where $\{\mathbf{H}^{1},\mathbf{H}^{2},\cdots,\mathbf{H}^{L}\}$ denotes the embedding matrix of the each layer of the graph encoder, respectively.
+where $\{\mathbf{H}^{1},\mathbf{H}^{2},\cdots,\mathbf{H}^{L}\}$ denotes the embedding matrix of each layer of the graph encoder, respectively.
 $\Theta=(\theta^{1},\ldots,\theta^{L})$ is the collection of weights across the layers.
 Pre-training.
-As stated by prior studies (Yu et al., 2024d, 2025), all contrastive pre-training task on graphs (Liu et al., 2023; Velickovic et al., 2019; You et al., 2020) can be unified under the task template of similarity calculation.
+As prior studies (Yu et al., 2024c, 2025c) have suggested, mainstream contrastive pre-training tasks on graphs (Liu et al., 2023; Velickovic et al., 2019; You et al., 2020) can be unified under the task template of similarity calculation.
 Formally, the unified pre-training objective is defined as follows:
-(3) $\mathcal{L}(\Theta)=-\sum_{o\in\mathcal{T}_{\text{pre}}}\ln\frac{\sum_{a\in Pos_{o}}\exp(\text{sim}(\mathbf{h}_{a},\mathbf{h}_{o})/\tau)}{\sum_{b\in Neg_{o}}\exp(\text{sim}(\mathbf{h}_{b},\mathbf{h}_{o})/\tau)},$
-where $Pos_{o}$ and $Neg_{o}$ denote the sets of positive and negative samples for a target instance $o$, respectively.
+(3) $\mathcal{L}(\Theta)=-\sum_{o\in\mathcal{T}_{\text{pre}}}\ln\frac{\sum_{a\in\text{Pos}_{o}}\exp(\text{sim}(\mathbf{h}_{a},\mathbf{h}_{o})/\tau)}{\sum_{b\in\text{Neg}_{o}}\exp(\text{sim}(\mathbf{h}_{b},\mathbf{h}_{o})/\tau)},$
+where $\text{Pos}_{o}$ and $\text{Neg}_{o}$ denote the sets of positive and negative samples for a target instance $o$, respectively.
 $\mathbf{h}_{o}$ represents the embedding of the target instance, while $\mathbf{h}_{a}$ and $\mathbf{h}_{b}$ correspond to the embeddings of positive and negative samples.
-The hyperparameter $\tau$ controls the temperature scaling in the similarity computation.
-In our framework, we follow previous work (Liu et al., 2023; Yu et al., 2024d) by employing similarity calculation as the task template and using link prediction as the pre-training task.
+The hyperparameter $\tau$ controls the temperature scaling of the similarity function $\text{sim}(\cdot,\cdot)$.
+In our framework, we follow previous work (Liu et al., 2023; Yu et al., 2024c) by employing similarity calculation as the task template and using link prediction as the pre-training task.
 Problem definition.
-In this work, we explore chain-of-thought (CoT) prompt learning framework for text-free graphs.
+In this work, we explore a CoT prompt learning framework for text-free graphs.
 We focus on two widely used tasks in graph learning: node classification and graph classification, in few-shot scenarios.
@@
 In contrast, graph classification considers a collection of graphs $\mathcal{G}$, where each graph $G_{i}\in\mathcal{G}$ is assigned a class label $Y_{i}\in Y$.
-In the few-shot setting, only $m$ labeled examples per class are available (e.g., $m\leq 10$), a paradigm defined as $m$-shot classification (Liu et al., 2023; Yu et al., 2024d).
+In the few-shot setting, only $m$ labeled examples per class are available (e.g., $m\leq 10$), which is referred to as $m$-shot classification (Liu et al., 2023; Yu et al., 2024c).
 4. Chain-of-Thought Graph Prompt Learning
-In this section, we propose our model, GCoT, starting with an overview of its framework. Then we detail its core components and conclude with a complexity analysis of the algorithm.
+In this section, we present our approach, GCoT, starting with an overview of its framework. Then we detail its core components and conclude with a complexity analysis of the algorithm.
 4.1. Overall Framework
-We illustrate the overall framework of GCoT in Fig.
-2, consisting of two key stages: pre-training and CoT prompting.
-First, we pre-train a graph encoder as shown in Fig.
-2(a).
-Details of pre-training are provided in Sect.
-3.
-Second, given a pre-trained graph encoder, to guide the model take additional inference step before finalizing predictions, we propose CoT prompting, as shown in Fig.
+We illustrate the overall framework of GCoT in Fig. 2, which consists of two phases: pre-training and CoT prompting.
+First, we pre-train a graph encoder as shown in Fig. 2(a). Details of the pre-training phase are provided in Sect. 3.
+Second, given a pre-trained graph encoder, to guide the model with additional inference steps before finalizing its predictions, we propose CoT prompting, as illustrated in Fig.
 2(c).
 Specifically, we design an inference step with three substages: prompt-based inference, thought construction, and thought-conditioned prompt learning.
-We first feed the prompt modified query graph into the pre-trained graph encoder.
-Then we construct a thought by fusing embeddings of all hidden layers of the pre-trained graph encoder.
-Conditioned on the thought from the previous step, we employ a condition-net to generate a series of node-specific conditional prompts that capture individualized learning patterns for nodes in a fine-grained and parameter-efficient manner, as shown in in Fig.
-2(b).
-We repeat the inference steps until obtain a final answer, as shown in Fig.
-2(c).
-Moreover, to enhance alignment between downstream tasks and the pre-training objective, we incorporate an initial prompt to adjust features or embeddings, same as prior graph prompting methods (Liu et al., 2023; Fang et al., 2024).
+Specifically, in each inference step, we first feed the prompt and query graph into the pre-trained graph encoder.
+Then, we construct a “thought” by fusing embeddings of all hidden layers of the pre-trained graph encoder.
+Lastly, as detailed in Fig.
+2(b), conditioned on the constructed thought, we employ a condition-net to generate a series of node-specific prompts.
+These conditional prompts capture individualized patterns for nodes in a fine-grained and parameter-efficient manner.
+We repeat the inference step one or more times before making the final prediction for the downstream task.
+It is worth noting that, to better align the downstream task with the pre-training task, we still incorporate a standard prompt similar to those used in prior graph prompting methods (Liu et al., 2023; Fang et al., 2024).
 4.2. Chain-of-Thought Prompting
-To bridge the gap between the objectives between pre-training and downstream tasks, we adhere to previous graph prompting methods, leveraging an initial prompt to modify node features (Sun et al., 2022; Fang et al., 2024) or output embeddings (Liu et al., 2023; Yu et al., 2025), as shown in Fig.
-2(c).
-Note that our framework is fully compatible with any standard graph prompting techniques and can further enhance their performance, as demonstrated in Sect.
-5.5.
-In our experiments, we employ an attention mechanism to generate prompts that modify the output embedding of the final step, as detailed in the Prompt tuning section.
-Prompt-based inference For the $k$-th inference step, we first feed the prompt modified query graph into the pre-trained encoder:
-(4) $\{\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}\}=\textsc{GraphEncoder}(\mathbf{X}_{k},G;\Theta_{0})$
-where $\Theta_{0}$ is the pre-trained parameters in the graph encoder, $\mathbf{X}_{k}$ is the modified feature by prompts, which will be introduced in the Thought conditioned prompt learning section.
-Thought construction To leverage the hierarchical knowledge across multiple layers of the graph encoder, we construct the thought by fusing the hidden embeddings from each layer of the pre-trained graph encoder as follows:
-(5) $\mathbf{T}_{k}=\mathtt{Fuse}(\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}),$
-where $\mathbf{H}_{k}^{l}$ denotes the hidden embedding from the $l$-th layer during the $k$-th inference step.
-The $\mathtt{Fuse}(\cdot)$ function can be implemented in various type.
-In our experiments, we employ weighted summation as the fusion function:
+Next, we introduce the details of each inference step and the standard prompt learning employed in our framework.
+4.2.1. Inference step
+Each inference step consists of three substages, as described in the following.
+Prompt-based inference. In the $k$-th inference step, we first feed the query graph $G=(V,E)$, along with its prompt-modified feature matrix, into the pre-trained encoder:
+(4) $\{\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}\}=\textsc{GraphEncoder}(\mathbf{X}_{k},G;\Theta_{0}),$
+where $\Theta_{0}$ denotes the frozen pre-trained weights in the graph encoder, and $\mathbf{X}_{k}$ is the node feature matrix modified by the prompts generated in the previous (i.e., $k-1$-th) step.
+The generation of the prompts and the modification of the feature matrix are elaborated later in thought-conditioned prompt learning.
+Note that for the first step ($k=1$), the original feature matrix is used without modification, i.e., $\mathbf{X}_{1}=\mathbf{X}$.
+Thought construction.
+To leverage the hierarchical knowledge across multiple layers of the graph encoder, we construct a “thought” by fusing the hidden embeddings from each layer of the pre-trained graph encoder as follows:
+(5) $\mathbf{T}_{k}=\textsc{Fuse}(\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}),$
+where each $\mathbf{H}_{k}^{l}$ denotes the hidden embedding from the $l$-th layer during the $k$-th inference step.
+The $\textsc{Fuse}(\cdot)$ function can be implemented in various ways.
+For example, if all hidden layers of the graph encoder share the same dimensionality, we can adopt a simple weighted summation [^1] This is the case in our experiments.
+If the dimensionalities differ, linear layers can be potentially used to project them to a common dimension., as follows.
 (6) $\mathbf{T}_{k}=w^{1}\cdot\mathbf{H}_{k}^{1}+w^{2}\cdot\mathbf{H}_{k}^{2}+\cdots+w^{L}\cdot\mathbf{H}_{k}^{L},$
-where $w^{1},w^{2},\cdots,w^{L}$ are learnable parameters. The thought $\mathbf{T}_{k}$ is then used to guide the the next inference step.
-Thought conditioned prompt learning The thought generated in the previous inference step stores hierarchical structural knowledge, which guides the learning in the subsequent step.
-Since different nodes may exhibit distinct characteristics with respect to the downstream task, rather than learn all node’s representation in the same manner, the pre-trained model should be steered to learn node-specific patterns.
-To this end, we propose leveraging a condition-net (Zhou et al., 2022; Yu et al., 2024b, 2025) to generate node-specific prompts, guiding node-specific adaptation.
-Specifically, conditioned on the thought from the previous step, $\mathbf{T}_{k}$, the condition-net generates a series of thought prompts as follows:
-(7) $\mathbf{P}_{k}=\mathtt{CondNet}(\mathbf{T}_{k};\phi),$
-where $\mathtt{CondNet}$ is the condition-net parameterized by $\phi$.
-Note that condition-net is a lightweight hypernetwork (Ha et al., 2022), where our condition-net $\mathtt{CondNet}$ functions as an auxiliary network to generate the prompts conditioned on the thoughts parameter-efficiently.
-In our implementation, we utilize a simple multi-layer perceptron (MLP) with a bottleneck structure for improved efficiency (Wu and Lee, 2018).
-The $i$-th row of $\mathbf{P}_{k}$ represents a unique node-specific prompt vector $\mathbf{p}_{i,k}$ for node $v_{i}$, which is derived from the thought $\mathbf{T}_{k}$.
-This prompt vector then guides the next inference step by modifying the node features as follows:
+where $W=\{w^{l}\in\mathbb{R}:1\leq l\leq L\}$ are learnable parameters.
+The resulting thought, $\mathbf{T}_{k}$, captures the current working state of the graph encoder, storing multi-layer structural knowledge.
+In particular, the $i$-th row of $\mathbf{T}_{k}$ reflects the thought pertaining specifically to node $v_{i}$.
+Thought-conditioned prompt learning.
+The thought constructed in Eq.
+(6) is then used for prompt generation, in order to further guide the next inference step.
+Moreover, since different nodes may exhibit distinct characteristics with respect to the downstream task, it is beneficial to adapt the pre-trained model to node-specific patterns.
+Thus, rather than employing a single prompt for all nodes, we propose to leverage a conditional network (condition-net) (Zhou et al., 2022; Yu et al., 2025b, c) to generate node-specific prompts and facilitate node-specific adaptation.
+Specifically, conditioned on the current thought $\mathbf{T}_{k}$, the condition-net generates a series of prompts, collectively represented as $\mathbf{P}_{k}\in\mathbb{R}^{|V|\times d}$, as follows.
+(7) $\mathbf{P}_{k}=\textsc{CondNet}(\mathbf{T}_{k};\phi),$
+where CondNet is the condition-net parameterized by $\phi$.
+The condition-net can be viewed as a hypernetwork (Ha et al., 2022)—a lightweight auxiliary network such as a multi-layer perceptron (MLP)—that generates a distinct prompt for each node from the thought, yet avoids parameterizing a separate prompt vector for each node.
+In particular, the $i$-th row of $\mathbf{P}_{k}$ represents a unique node-specific prompt vector $\mathbf{p}_{k,i}$ for node $v_{i}\in V$.
+Subsequently, the prompt vectors are used to modify the node features of the query graph for prompt-based inference in the next (i.e., $k+1$-th) step, as follows.
 (8) $\mathbf{X}_{k+1}=\mathbf{P}_{k}\odot\mathbf{X},$
-where $\odot$ denotes element-wise multiplication, and $\mathbf{X}_{k+1}$ represents the input to the pre-trained graph encoder in the $k+1$ step.
-After repeating $K$ inference steps, the final step outputs the nodes embeddings $\mathbf{H}_{K}$.
-Prompt tuning For initial prompts, all standard graph prompting methods can be applied within our framework.
-In our experiments, we adhere to GPF+ (Fang et al., 2024) to generate initial prompts.
-Specifically, we train $N$ bias prompts $\{\mathbf{p}_{\text{bias}}^{1},\mathbf{p}_{\text{bias}}^{2},\ldots,\mathbf{p}_{\text{bias}}^{N}\}$, and leverage attention-based aggregation to generate node-specific prompts.
-However, unlike GPF+, which uses these prompts to modify graph features, we use them to modify the final embeddings.
-Specifically, the task prompt for node $i$ is computed as:
-(9) $\mathbf{p}_{\text{init}}^{i}=\sum_{j=1}^{N}\alpha_{i,j}\mathbf{p}_{\text{bias}}^{j},\quad\text{where}\quad\alpha_{i,j}=\frac{\exp\big{(}\mathbf{a}^{j}\mathbf{h}_{K}^{i}\big{)}}{\sum_{l=1}^{N}\exp\big{(}\mathbf{a}^{l}\mathbf{h}_{K}^{i}\big{)}}.$
-Here, $\mathbf{h}_{K}^{i}$ denotes the $i$-th row of the final embedding matrix $\mathbf{H}_{K}$, which serves as the embedding for node $i$, and $\mathbf{A}=\{\mathbf{a}^{1},\mathbf{a}^{2},\ldots,\mathbf{a}^{N}\}$ are $N$ learnable linear projection vectors.
-The generated initial prompts form the matrix $\mathbf{P}_{\text{init}}$, which is then used to modify the final output embeddings as follows:
-(10) $\mathbf{\tilde{H}}=\mathbf{P}_{\text{init}}\odot\mathbf{H}_{K}.$
-Note that for clarity and to maintain a unified representation of the various task prompt mechanisms, we use $\mathbf{P}_{\text{init}}$ to denote the trainable parameters associated with these mechanisms throughout the remainder of this paper.
-Finally, for a given task with a labeled training set
-$\mathcal{D}_{t}=\{(x_{1},y_{1}),(x_{2},y_{2}),\dots\},$
-where each $x_{i}$ represents either a node or a graph and $y_{i}\in Y$ denotes its corresponding class label, the downstream loss function is defined as:
-(11) $\mathcal{L}_{\text{down}}(\phi,\mathbf{P}_{\text{init}})=-\sum_{(x_{i},y_{i})\in\mathcal{D}_{t}}\ln\frac{\exp\left(\text{sim}\left(\mathbf{\tilde{h}}_{x_{i}},\mathbf{\bar{h}}_{y_{i}}\right)/\tau\right)}{\sum_{c\in Y}\exp\left(\text{sim}\left(\mathbf{\tilde{h}}_{x_{i}},\mathbf{\bar{h}}_{c}/\right)\tau\right)},$
+where $\odot$ denotes element-wise multiplication, and $\mathbf{X}_{k+1}$ represents the input to the pre-trained graph encoder in the $k+1$-th step.
+4.2.2. Standard prompt learning
+To align the objectives of pre-training and downstream tasks, we also leverage a standard prompt following prior work, which typically modifies node features (Sun et al., 2022; Fang et al., 2024; Yu et al., 2024c) or embeddings (Liu et al., 2023; Yu et al., 2024c).
+In particular, we adopt GPF+ (Fang et al., 2024) to generate the standard prompts.
+However, we emphasize that our framework is compatible with any standard graph prompting technique.
+As demonstrated in Sect.
+5.6, our CoT-style prompting can be combined with several prevailing graph prompting approaches to further enhance their performance.
+Specifically, we train $N$ bias prompts $\{\mathbf{p}_{\text{bias}}^{1},\ldots,\mathbf{p}_{\text{bias}}^{N}\}$ and leverage attention-based aggregation to generate node-specific prompts.
+While GPF+ (Fang et al., 2024) applies these prompts to graph features $\mathbf{X}$, we use them to modify the output embeddings $\mathbf{H}_{K}$ after $K$ inference steps.
+Concretely, the standard prompt for node $v_{i}$ is computed as
+(9) $\mathbf{p}_{\text{std},i}=\sum_{j=1}^{N}\alpha_{i,j}\mathbf{p}_{\text{bias}}^{j},\quad\text{where}\ \alpha_{i,j}=\frac{\exp\big{(}\mathbf{a}^{j}\mathbf{h}_{K,i}\big{)}}{\sum_{n=1}^{N}\exp\big{(}\mathbf{a}^{n}\mathbf{h}_{K,i}\big{)}}.$
+Here, $\mathbf{h}_{K,i}$ denotes the $i$-th row of the output embedding matrix $\mathbf{H}_{K}$, i.e., the embedding vector of node $v_{i}$ after $K$ inference steps.
+$\{\mathbf{a}^{1},\mathbf{a}^{2},\ldots,\mathbf{a}^{N}\}$ are $N$ learnable linear projection vectors.
+The standard prompts for all nodes are stacked to form the matrix $\mathbf{P}_{\text{std}}$, which is then used to modify the output embeddings $\mathbf{H}_{K}$:
+(10) $\mathbf{\tilde{H}}=\mathbf{P}_{\text{std}}\odot\mathbf{H}_{K}.$
+We call $\mathbf{\tilde{H}}$ the answer matrix, since it is used by downstream tasks to produce final predictions (or answers) in Sect. 4.3.
+Note that for clarity and to maintain a unified representation of various standard prompt mechanisms, we use $\mathbf{P}_{\text{std}}$ to denote the trainable parameters associated with these mechanisms throughout the remainder of this paper.
+4.2.3. Summary
+In GCoT, we adopt the following mechanisms, which collectively ensure the effectiveness of GCoT.
+(1) We follow standard graph prompt learning methods (Liu et al., 2023; Fang et al., 2024) to align pre-training and downstream tasks, ensuring that GCoT can efficiently adapt to different downstream tasks even in few-shot settings.
+(2) The CoT-style prompting performs multiple inference steps, allowing iterative refinement of its prediction for a given downstream task.
+(3) The thoughts in GCoT fuse hierarchical topological knowledge from graphs, enabling the capture of fine-grained structural information.
+(4) Conditioned on the thought, GCoT generates a series of node-specific prompts that reflect individualized node characteristics in a parameter-efficient manner.
+4.3. Prompt Tuning
+Consider a downstream task with a labeled training set
+$\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\dots\},$
+where each $x_{i}$ represents either a node or a graph and $y_{i}\in Y$ denotes its corresponding class label.
+Subsequently, the loss function of the task is defined as $\mathcal{L}_{\text{down}}(\mathcal{D};W,\phi,\mathbf{P}_{\text{std}})=$
+(11) $-\sum_{(x_{i},y_{i})\in\mathcal{D}}\ln\frac{\exp\left(\text{sim}\left(\mathbf{\tilde{h}}_{x_{i}},\mathbf{\tilde{h}}_{y_{i}}\right)/\tau\right)}{\sum_{c\in Y}\exp\left(\text{sim}\left(\mathbf{\tilde{h}}_{x_{i}},\mathbf{\tilde{h}}_{c}\right)/\tau\right)},$
 where $\mathbf{\tilde{h}}_{x_{i}}$ represents the final embedding of a node $v$ or a graph $G$.
-Specifically, for node classification, $\mathbf{\tilde{h}}_{v}$ corresponds to a row in the answer matrix $\mathbf{\tilde{H}}$, while for graph classification, we compute the graph embedding as
-$\mathbf{\tilde{h}}_{G}=\sum_{v\in V}\mathbf{\tilde{h}}_{v},$
-incorporating an additional graph readout step.
-The prototype embedding for each class $c$, denoted as $\mathbf{\bar{h}}_{c}$, is obtained by averaging the embeddings of all labeled nodes or graphs belonging to that class.
-During prompt tuning, only the task prompts and lightweight parameters of the condition network ($\phi$) are updated, while the pre-trained GNN weights remain frozen.
-This parameter-efficient design makes our approach well-suited for few-shot learning, where the training set $\mathcal{D}_{t}$ contains only a limited number of labeled examples.
+Specifically, for node classification, $\mathbf{\tilde{h}}_{v}$ corresponds to a row in the answer matrix $\mathbf{\tilde{H}}$; for graph classification, we apply a readout operation and compute the graph embedding as $\mathbf{\tilde{h}}_{G}=\sum_{v\in V}\mathbf{\tilde{h}}_{v}$.
+Lastly, $\mathbf{\tilde{h}}_{c}$ denotes the prototype embedding for class $c$, which is obtained by averaging the embeddings of all labeled nodes or graphs belonging to that class.
+During prompt tuning, only the weights of thought construction ($W$) and the condition-net ($\phi$), as well as the standard prompts ($\mathbf{P}_{\text{std}}$), are updated, while the pre-trained weights of the graph encoder remain frozen.
+This parameter-efficient design makes our approach well-suited for few-shot learning, where the training set $\mathcal{D}$ contains only a few labeled examples.
 4.4. Algorithm and Complexity Analysis
-Algorithm.
-We detail the main steps for Chain-of-Thought graph prompting in Algorithm 1.
-In lines 3–12, we iterate through $K$ inference steps during the downstream adaptation phase while keeping the pre-trained weights $\Theta_{0}$ frozen.
-Specifically, in lines 4–7, we generate the thought for the $k$-th inference step by fusing the output embeddings from each layer of the pre-trained graph encoder.
-In lines 8–12, we leverage this thought to generate node-specific prompts that guide the subsequent inference step to capture node-specific patterns.
-In lines 13–14, we employ a standard task prompt to modify the output embedding of the last inference step.
-Note that, based on the task prompt mechanism described in the original paper, these prompts can also be applied to modify the input features at the first inference step.
-Finally, in lines 15–17, we update the embeddings for the prototypical nodes/graphs based on the few-shot labeled data.
 Algorithm 1 Chain-of-Thought Graph Prompt Learning
-1:Pre-trained graph encoder with parameters $\Theta_{0}$.
-2:Optimized parameters $\phi$ of condition-net, and task prompt $\mathbf{P}_{\text{init}}$.
-3:while not converged do
-4:    $\phi_{i}\leftarrow$ initialization
+1:Pre-trained graph encoder with parameters $\Theta_{0}$, labeled data $\mathcal{D}$ for a given downstream task
+2:Optimized parameters $W,\phi,\mathbf{P}_{\text{std}}$.
+3:$W,\phi,\mathbf{P}_{\text{std}}\leftarrow$ initialization
+4:while not converged do
 5:    $\mathbf{X}_{1}\leftarrow\mathbf{X}$
-6:    /* 1 to $K-1$ inference steps*/
+6:    /* First $K-1$ inference steps*/
 7:    while inference step $1\leq k<K$ do
-8:         /* Encoding by the pre-trained graph encoder */
-9:         $\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}\leftarrow\textsc{GraphEncoder}(G,\mathbf{X}_{k};\Theta_{0})$
+8:         /* Prompt-based inference */
+9:         $\{\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L}\}\leftarrow\textsc{GraphEncoder}(G,\mathbf{X}_{k};\Theta_{0})$
 10:         /* Thought construction */
 11:         $\mathbf{T}_{k}\leftarrow\mathtt{Fuse}(\mathbf{H}_{k}^{1},\mathbf{H}_{k}^{2},\cdots,\mathbf{H}_{k}^{L})$
-12:         /* Thought prompting */
-13:         /* Generate thought prompts by Eq. 7 */
-14:         $\mathbf{P}_{k}\leftarrow\mathtt{CondNet}(\mathbf{T}_{k};\phi)$
-15:         /* Thought prompts modification by Eq. 8 */
-16:         $\mathbf{X}_{k+1}\leftarrow\mathbf{P}_{k}\odot\mathbf{X}$
-17:    /* $K$-th inference step */
-18:    $\{\mathbf{H}_{K}^{1},\mathbf{H}_{K}^{2},\cdots,\mathbf{H}_{K}^{L}\}\leftarrow\textsc{GraphEncoder}(G,\mathbf{X}_{K};\Theta_{0})$
-19:    $\mathbf{H}_{K}\leftarrow\mathbf{H}_{K}^{L}$
-20:    /* Initial prompt modification by Eq. 10*/
-21:    $\mathbf{\tilde{H}}\leftarrow\mathbf{P}_{\text{init}}\odot\mathbf{H}_{K}$
-22:    /* Update prototypical instance */
-23:    for each class $c$ do
-24:         ${\mathbf{\bar{h}}}_{c}\leftarrow\textsc{Average}(\mathbf{\tilde{h}}_{x}$: instance $x$ belongs to class $c$)
-25:    /* Optimizing the parameters in condition-net */
-26:    Calculate $\mathcal{L}_{\text{down}}(\phi,\mathbf{P}_{\text{init}})$ by Eq. (11)
-27:    Update $\phi$ by backpropagating $\mathcal{L}_{\text{down}}(\phi,\mathbf{P}_{\text{init}})$
-28:return $\{\phi,\mathbf{P}_{\text{task}}\}$
+12:         /* Thought-conditioned prompt generation by Eq. (7) */
+13:         $\mathbf{P}_{k}\leftarrow\mathtt{CondNet}(\mathbf{T}_{k};\phi)$
+14:         /* Feature modification by Eq. (8) */
+15:         $\mathbf{X}_{k+1}\leftarrow\mathbf{P}_{k}\odot\mathbf{X}$
+16:    /* Last (i.e., $K$-th) inference step */
+17:    $\{\mathbf{H}_{K}^{1},\mathbf{H}_{K}^{2},\cdots,\mathbf{H}_{K}^{L}\}\leftarrow\textsc{GraphEncoder}(G,\mathbf{X}_{K};\Theta_{0})$
+18:    $\mathbf{H}_{K}\leftarrow\mathbf{H}_{K}^{L}$
+19:    /* Standard prompt modification by Eq. (10) */
+20:    $\mathbf{\tilde{H}}\leftarrow\mathbf{P}_{\text{std}}\odot\mathbf{H}_{K}$
+21:    /* Update prototypical instance */
+22:    for each class $c$ do
+23:         ${\mathbf{\tilde{h}}}_{c}\leftarrow\textsc{Mean}(\{\mathbf{\tilde{h}}_{x}$: $x$ is any instance in class $c$ from $\mathcal{D}$})
+24:    /* Optimizing the parameters */
+25:    Calculate $\mathcal{L}_{\text{down}}(\mathcal{D};W,\phi,\mathbf{P}_{\text{std}})$ by Eq. (11)
+26:    Update $W,\phi,\mathbf{P}_{\text{std}}$ by backpropagating $\mathcal{L}_{\text{down}}(\mathcal{D};W,\phi,\mathbf{P}_{\text{std}})$
+27:return $\{W,\phi,\mathbf{P}_{\text{std}}\}$
+Algorithm.
+We outline the main steps for GCoT in Algorithm 1.
+In lines 3–13, we iterate through the first $K-1$ inference steps for a downstream task while keeping the pre-trained weights $\Theta_{0}$ frozen.
+Specifically, in lines 8–9, we generate the thought for the $k$-th inference step by fusing the embeddings from each layer of the pre-trained graph encoder.
+In lines 10–11, we leverage this thought to generate node-specific prompts that guide the next inference step.
+In lines 14–18, we perform the last inference step, employing a standard prompt to modify the output embedding and obtaining the answer matrix.
+Finally, in lines 19–24, we calculate and backpropagate the downstream loss to update the parameters $W,\phi,\mathbf{P}_{\text{std}}$.
 Complexity analysis.
-For a downstream graph $G$, we perform $K$ iterative inference steps.
-Each step comprises two key components: (1) thought generation using a pre-trained GNN (executed for $K$ iterations) and (2) thought-based conditional prompt learning (executed for $K-1$ iterations).
-The computational complexity of the first component is largely determined by the GNN architecture.
+For a downstream graph $G$, we perform $K$ inference steps.
+In each step, we first conduct prompt-based inference, the complexity of which is determined by the architecture of the graph encoder.
 In a standard GNN, each node aggregates information from up to $D$ neighboring nodes per layer.
-Consequently, computing node embeddings over $L$ layers incurs a complexity of $O(D^{L}\cdot|V|)$, where $|V|$ denotes the number of nodes.
-Aggregating the outputs from all $L$ layers adds an extra complexity of $O(L\cdot|V|)$.
-Since the thought generation process is repeated $K$ times, its overall complexity is:$O(K\cdot(D^{L}+L\cdot|V|))$.
-The second component, thought-based conditional prompt learning, refines the learning process by leveraging the thought representation from the previous step.
-This component consists of two stages: prompt generation and prompt tuning.
-In the prompt generation stage, the thought is fed into the condition-net, with a complexity of $O(|V|)$.
-In the prompt tuning stage, each node is adjusted via a prompt vector, also incurring a complexity of $O(|V|)$.
-Since this process is executed for $K-1$ iterations, the total complexity for this phase is: $O(2(K-1)\cdot|V|)$.
-Additionally, we employ a standard task prompt to modify node features or embeddings.
+Consequently, computing node embeddings over $L$ layers results in a complexity of $O(D^{L}|V|)$, where $|V|$ denotes the number of nodes.
+This is followed by thought generation, which fuses embeddings from all $L$ layers, introducing an additional complexity of $O(L|V|)$.
+The generated thought is subsequently used for thought-conditioned prompt learning, with a complexity of $O(|V|)$.
+Therefore, the computational complexity for the $K$ inference steps is $O(K(D^{L}+L)|V|)$.
+Additionally, we employ a standard prompt to modify node features or embeddings.
 The complexity of this process depends on the specific prompting mechanism.
 Here, we assign it a complexity of $O(|V|)$, which is common among graph prompting methods (Liu et al., 2023; Fang et al., 2024).
-In summary, the overall computational complexity of GCoT is: $O((D^{L}+L+2(K-1)+1)\cdot|V|)$.
-4.4. Why GCoT Works
-In GCoT, we adapt the following mechanism.
-(1) All three approaches leverage a universal task template to unify pre-training and downstream tasks, ensuring that GCoT can efficiently adapt to different downstream tasks, particularly in few-shot settings.
-(2) The chain-of-thought mechanism enables GCoT to perform multiple inference steps, thereby refining the final answer more effectively.
-(3) The thoughts in GCoT fuse hierarchical topological knowledge from graphs, allowing it to better capture structural information.
-(4) Conditioned on the thought from the previous step, GCoT generates a series of node-specific prompts that efficiently capture fine-grained node characteristics, thus effectively refining answer step by step.
-The combination of above mechanisms ensures the effectiveness of GCoT.
+Thus, the overall complexity of GCoT is $O(K(D^{L}+L)|V|)$, or more simply $O(KD^{L}|V|)$, since it is typical that $L\ll D^{L}$.
 5. Experiments
@@
 Citeseer 1 - 3,327 4,732 3,703 6 N
-Pubmed 1 - 19717 88648 500 3 N
-Photo 1 - 7650 238162 745 8 N
+Pubmed 1 - 19,717 88,648 500 3 N
+Photo 1 - 7,650 238,162 745 8 N
 MUTAG 188 2 17.9 18.9 7 - G
@@
 PROTEINS 1,113 2 39.1 72.8 4 3 G
-indicates the downstream task category for each dataset: ”N” corresponds to node classification, while ”G” signifies graph classification.
-Table 2. Accuracy (%) evaluation on node and graph classification.
+This column indicates the type of downstream task conducted for each dataset: “N” denotes node classification, while “G” denotes graph classification.
+Table 2. Accuracy (%) evaluation of node and graph classification.
 Methods Node classification Graph classification
@@
 GAT 31.00 ± 16.22 27.71 ±   8.74 50.02 ±   8.88 51.79 ± 12.85 37.33 ± 10.81 50.58 ± 26.16 46.55 ± 16.57 40.39 ±   0.04
-DGI/InfoGraph 54.11±9.60 45.00±9.19 57.98±11.36 64.48±10.04 53.17±17.29 53.82±14.19 49.33±15.11 52.51±10.29
-GraphCL 51.96±9.43 43.12±9.61 61.18±11.60 63.93±10.21 54.92±17.09 53.81±14.21 49.73±14.66 53.81±8.97
-ProG 51.35±13.92 35.08±8.63 64.07±14.74 68.85±9.99 51.99±4.50 53.45±15.01 53.52±11.97 52.73±6.57
-GPF 57.35±13.87 43.04±8.85 57.61±11.26 65.26±9.89 56.55±13.95 54.16±14.07 48.65±13.96 53.05±7.62
-GPF+ 64.91±15.49 42.28±8.71 57.85±10.86 69.80±10.84 56.81±12.93 55.24±13.29 50.83±19.74 54.58±8.70
-GraphPrompt 54.56±12.56 46.23±8.70 64.85±10.72 68.99±9.28 55.44±12.56 54.34±14.77 54.59±10.52 53.80±7.93
-GCoT 66.09±15.17 47.77±10.93 70.57±17.10 77.51±11.76 58.75±15.42 56.26±15.52 58.03±23.44 56.24±8.60
+DGI/InfoGraph 54.11 ±   9.60 45.00 ±   9.19 47.46 ± 12.19 58.89 ± 10.97 53.17 ± 17.29 53.82 ± 14.19 49.33 ± 15.11 52.51 ± 10.29
+GraphCL 51.96 ±   9.43 43.12 ±   9.61 46.80 ±   9.04 57.78 ± 11.31 54.92 ± 17.09 53.81 ± 14.21 49.73 ± 14.66 53.81 ±   8.97
+ProG 50.59 ± 14.64 43.17 ±   8.49 63.07 ± 11.96 66.50 ±   9.46 51.99 ±   4.50 53.45 ± 15.01 53.52 ± 11.97 52.73 ±   6.57
+GPF 57.60 ± 13.88 43.11 ±   8.80 55.63 ± 10.96 65.29 ± 10.07 56.55 ± 13.95 54.16 ± 14.07 48.65 ± 13.96 53.05 ±   7.62
+GPF+ 57.42 ± 13.87 43.28 ±   8.82 57.16 ± 10.99 65.07 ± 10.01 56.81 ± 12.93 55.24 ± 13.29 50.83 ± 19.74 54.58 ±   8.70
+GraphPrompt 54.25 ±   9.38 45.34 ± 10.53 63.11 ± 10.01 66.62 ±   9.90 55.44 ± 12.56 54.34 ± 14.77 54.59 ± 10.52 53.80 ±   7.93
+GCoT 59.67 ± 15.51 46.21 ±  8.78 64.43 ±  9.96 67.16 ± 10.46 58.75 ± 15.42 56.26 ± 15.52 58.03 ± 23.44 56.24 ±  8.60
 Best results are bolded and runner-up results are underlined.
@@
 In these datasets, nodes represent academic papers and edges denote citation relationships between them.
-Photo (Shchur et al., 2018) is an e-commerce co-purchase network derived from Amazon’s photography-related product category, where nodes correspond to products and edges indicate frequently co-purchased items.
+Photo (Shchur et al., 2018) is an e-commerce co-purchase network derived from Amazon’s photography-related product categories, where nodes correspond to products and edges indicate frequently co-purchased items.
 PROTEINS (Borgwardt et al., 2005) is a dataset of protein structures.
 In each graph, nodes correspond to secondary structures, and edges capture spatial or sequential relationships within the amino acid sequence.
-MUTAG (Rossi and Ahmed, 2015), BZR (Rossi and Ahmed, 2015), and COX2 (Rossi and Ahmed, 2015) are molecular graph datasets, representing nitroaromatic compounds, ligands associated with benzodiazepine receptors, and COX2 contains molecular structures related to cyclooxygenase-2 inhibitors, respectively.
-A detailed summary of these datasets is presented in Table 1, with further descriptions in Appendix A.
+MUTAG (Rossi and Ahmed, 2015), BZR (Rossi and Ahmed, 2015), and COX2 (Rossi and Ahmed, 2015) are molecular graph datasets, representing nitroaromatic compounds, ligands associated with benzodiazepine receptors, and molecular structures related to cyclooxygenase-2 inhibitors, respectively.
+These datasets are summarized in Table 1, with further descriptions in Appendix A.
 Baselines.
-We compare GCoT with state-of-the-art methods across three categories: (1) Supervised GNNs: GCN (Kipf and Welling, 2017) and GAT (Veličković et al., 2018) are trained directly on downstream labels in a fully supervised manner, without any pre-training.
+We compare GCoT with state-of-the-art methods across three categories.
+(1) Supervised GNNs: GCN (Kipf and Welling, 2017) and GAT (Veličković et al., 2018) are trained directly on downstream labels in a supervised manner, without any pre-training.
 (2) Graph pre-training models: DGI/InfoGraph [^2] DGI is originally designed for node-level tasks, while InfoGraph extends it to graph-level classification.
@@
 (Veličković et al., 2018; Sun et al., 2020) and GraphCL (You et al., 2020) adopt a “pre-train, fine-tune” strategy.
-These models first perform self-supervised pre-training using unlabeled graphs and are later fine-tuned for downstream tasks, where a classifier is trained with few-shot labels while keeping the pre-trained encoder frozen in single step.
+They first perform self-supervised pre-training using unlabeled graphs and are later fine-tuned for downstream tasks, where a classifier is trained with few-shot labels while keeping the pre-trained encoder frozen.
 (3) Graph prompt learning models: ProG (Sun et al., 2023), GPF (Fang et al., 2024), GPF+ (Fang et al., 2024), and GraphPrompt (Liu et al., 2023) employ self-supervised pre-training followed by prompt tuning.
-Unlike the fine-tuning methods, these methods leverage a unified task template, and train task-specific prompts in single step for downstream adaptation.
-Further descriptions of these baselines and implementation details are shown in Appendix B &  C, respectively.
+Unlike the fine-tuning methods, these methods leverage a unified task template, and train task-specific prompts for downstream adaptation.
+Further descriptions of these baselines are presented in Appendix B.
 Downstream tasks and evaluation.
 We perform experiments on two downstream tasks: node classification and graph classification.
-Both tasks follow an $m$-shot classification setup, where for each class, we randomly select $m$ instances (nodes or graphs) as labeled examples.
+Both types of task follow an $m$-shot classification setup, where for each class, we randomly select $m$ instances (nodes or graphs) as labeled examples.
+The remaining instances are treated as the test set.
 In our main results, we set $m=1$ for both node and graph classification tasks.
-Additionally, to examine the robustness of our method, we vary $m$ within the range $1\leq m\leq 10$, allowing us to analyze performance under different few-shot scenarios.
-We construct 100 independent $m$-shot tasks for each classification setting through repeated sampling.
+Additionally, to examine the robustness of our method, we vary $m$ within the range $[1,10]$, allowing us to analyze the performance under different few-shot scenarios.
+We construct 100 independent $m$-shot tasks for each type of classification through repeated sampling.
 Each task is evaluated using five different random seeds, resulting in a total of 500 experimental runs per task type.
-We report both the mean and standard deviation across these results.
-We evaluate performance using accuracy, in line with previous studies (Wang et al., 2020; Liu et al., 2021a, 2023).
-5.2. Performance Evaluation
-We first evaluate one-shot classification tasks. Then, we examine the effect of increasing the number of shots on model performance.
-One-shot performance.
-We present the results for one-shot node and graph classification tasks in Tables 2.
-We observe that: (1) GCoT consistently outperforms most baseline methods, surpassing the best competitor by margins of up to 11.05% in node classification and 6.03% in graph classification.
-These results underscore the advantage of incorporating multiple reasoning steps to better adapt to diverse tasks.
-(2) Standard graph prompt learning approaches—ProG, GPF, GPF+, and GraphPrompt—exhibit significantly weaker performance compared to GCoT.
-Their limitations arise from relying on a single-step inference process that lacks iterative refinement of the final answer.
-This contrast highlights the effectiveness of our chain-of-thought prompting, which enables step-by-step inference and captures individualized learning patterns for each node.
-Few-shot performance.
-To evaluate the impact of labeled data availability on GCoT’s performance, we vary the number of shots in both node and graph classification tasks.
-The results, shown in Fig.
-3, compare GCoT against several competitive baselines.
-First, GCoT consistently outperforms the baselines, particularly in low-shot scenarios (e.g., $m\leq 5$), where labeled data is scarce.
-Second, as the number of labeled samples increases (e.g., $m>5$), all methods generally exhibit improved performance, which aligns with expectations.
-Nevertheless, GCoT remains highly competitive, often achieving the best results.
-Note that on certain datasets, such as PROTEINS, performance across methods tends to show high variance.
-A possible reason is that this dataset exhibits greater variability in graph sizes compared to other datasets: The standard deviation of graph sizes in PROTEINS is 45.78, whereas other datasets fall within the range of 4.04 to 15.29.
-This may contribute to unstable performance.
-Despite this, GCoT demonstrates greater robustness than the competing methods.
-For the rest of the experiments, we focus on the one-shot classification setting.
-Figure 3. Impacts of shots on node and graph classification.
-5.3. Ablation Study and Visualization
-[…]
-[…]
-Table 3. Ablation study on the effects of key components.
-Methods Node classification Graph classification
-Cora Pubmed MUTAG PROTEINS
-[…]
-GCoT-L1 61.45±15.82 59.94±14.84 56.54±14.12 54.71±8.57
-GCoT-L2 66.95±16.07 63.05±15.13 57.68±13.84 54.77±8.81
-GCoT-L3 64.39±15.76 69.27±18.20 57.85±16.10 56.22±8.45
-GCoT 66.09±15.17 70.57±17.10 58.75±15.42 56.24±8.60
-Figure 4. Visualization of output embedding space of nodes. Different colors represents different classes.
-5.4. Heterophily Sensitivity
-To examine the robustness of GCoT on heterophilic graphs (Pei et al., 2020; Yu et al., 2025), we conduct one-shot node classification on heterophilic datasets (Wisconsin (Pei et al., 2020) and Squirrel (Rozemberczki et al., 2021)).
-We provide the detailed description of these datasets in Appendix D.
-As shown in Table 4, GCoT consistently outperforms other competing baselines, further validating its effectiveness.
-These results indicates that the iterative inference process in GCoT successfully generalizes across both homophilic and heterophilic graphs.
-Table 4. Accuracy (%) evaluation on heterophilic datasets.
-Methods Wisconsin Squirrel
-DGI 28.04 $\pm$ 6.47 20.00 $\pm$ 1.86
-GraphCL 29.85 $\pm$ 8.46 21.42 $\pm$ 2.22
-ProG 32.95 $\pm$ 11.42 20.61 $\pm$ 3.01
-GPF 30.29 $\pm$ 10.04 19.89 $\pm$ 3.19
-GPF+ 29.26 $\pm$ 9.61 20.93 $\pm$ 3.98
-GraphPrompt 29.96 $\pm$ 9.40 21.22 $\pm$ 1.80
-GCoT 33.04 $\pm$ 9.71 22.03 $\pm$ 4.94
-5.5. Flexibility of Graph Prompting Methods
-To evaluate the flexibility and robustness of GCoT, we evaluate its performance using various standard graph prompting methods as the task prompt.
-Specifically, we integrate ProG (Sun et al., 2023), GPF (Fang et al., 2024), GPF+ (Fang et al., 2024), and GraphPrompt (Liu et al., 2023) into our framework.
-The results for both node and graph classification on four datasets are presented in Table 5.
-Across nearly all cases, GCoT consistently outperforms its single-step prompting counterparts, regardless of the task prompt employed.
-These findings highlight the robustness of our approach and further confirm the advantage of chain-of-thought prompting over previous single-step prompting.
-Table 5. Evaluation of GCoT with different standard prompting methods.
-Prompting Thinking Node classification Graph classification
-Cora Pubmed MUTAG PROTEINS
-ProG - 51.35±13.92 64.07±14.74 51.99±14.50 52.73±6.57
-GCoT 59.62±14.20 72.19±14.92 58.42±15.39 55.82±8.79
-GPF - 57.35±13.87 57.61±11.26 56.55±13.95 53.05±7.62
-GCoT 60.28±15.69 66.24±12.61 57.66±13.11 56.98±8.30
-GPF+ - 64.91±15.49 57.85±10.86 56.81±12.93 54.58±8.70
-GCoT 62.71±15.18 65.97±11.09 58.54±13.29 55.45±9.20
-GraphPrompt - 54.56±12.56 64.85±10.72 55.44±12.56 53.80±7.93
-GCoT 55.83±13.69 72.73±16.18 58.28±15.47 55.32±8.43
-“-” refers to standard graph prompting without our CoT design.
-5.6. Hyperparameter Analysis for Condition-Net
-In our experiments, we implement the condition-net as a two-layer MLP with a bottleneck design.
-To examine its effect on performance, we vary the hidden dimension $s$ and present the results in Fig.
-5.
-Our findings reveal that $s=32$ generally yields competitive results for node classification, and $s=8$ for graph classification, leading us to adopt these settings in our experiments.
-A smaller $s$ may restrict the model’s representational capacity, limiting its effectiveness, whereas a larger $s$ introduces additional trainable parameters and increases the likelihood of overfitting in few-shot learning scenarios.
-Figure 5. Impact of hidden dimension $s$ in the condition-net.
-5.7. Effect of Varying Number of Inference Steps
-We further vary the number $K$ of inference steps to analyze their impact, with the results presented in Fig.
-6.
-We observe that for both node and graph classification tasks, $K=2$ generally yields optimal performance.
-Hence, we adopt $K=2$ in our experiments.
-For Proteins, $K=3$ achieves the best result, likely due to its inherent complexity as described in Sect.
-5.2.
-Figure 6. Impact of number of inference steps $K$.
-6. Conclusions
-In this paper, we propose GCoT, the first CoT prompting framework for graphs.
-We define an inference step with three substages: prompt-based inference, thought construction, and thought conditioned prompt learning.
-Specifically, we first feed the prompt modified query into the pre-trained encoder, and then construct a thought by fusing the hidden embeddings from each layer of the pre-trained graph encoder.
-To guide the subsequent inference step, we generate a series of prompts conditioned on the thought from the previous step.
-By repeating the above inference steps, GCoT obtain the answer.
-Finally, we conduct extensive experiments on eight public datasets, demonstrating that GCoT significantly outperforms a range of state-of-the-art baselines.
-Appendices
-A. Further Descriptions of Datasets
-In this section, we provide more comprehensive descriptions of the benchmark datasets used in our experiments, as summarized in Table 1, in the following.
-  • Cora [^2] https://github.com/shchur/gnn-benchmark/raw/master/data/npz/cora.npz (McCallum et al., 2000) is a citation network composed of 2,708 research papers in the field of computing, each classified into one of seven categories.
-The network consists of 5,429 citation links between papers.
-Each paper is represented by a binary word vector, where each entry indicates the presence or absence of a word from a predefined vocabulary of 1,433 unique terms.
-  • Citeseer [^3] https://github.com/shchur/gnn-benchmark/raw/master/data/npz/citeseer.npz (Sen et al., 2008) includes 3,312 computer science publications, categorized into six distinct classes, separate from those in Cora.
-The citation network contains 4,732 edges.
-Each document is encoded as a binary word vector that captures the presence or absence of words from a dictionary comprising 3,703 unique terms.
-  • PubMed [^4] https://github.com/shchur/gnn-benchmark/raw/master/data/npz/pubmed.npz (Sen et al., 2008) is a citation network of 19,717 biomedical articles related to diabetes, divided into three categories.
-The network includes 44,338 citation edges.
-Unlike Cora and Citeseer, each document is represented by a TF/IDF-weighted word vector derived from a dictionary of 500 unique terms.
-  • Photo [^5] https://github.com/shchur/gnn-benchmark/raw/master/data/npz/amazon_electronics_photo.npz (Shchur et al., 2018) consists of 7,487 photography-related products, each assigned to one of eight categories.
-The co-purchase network contains 119,043 edges, where connections indicate products frequently bought together.
-Each product is described by a feature vector extracted from metadata and customer reviews, with category labels corresponding to product types.
-  • MUTAG [^6] https://huggingface.co/datasets/graphs-datasets/MUTAG (Debnath et al., 1991) is a dataset of nitroaromatic compounds aimed at predicting their mutagenic effects on Salmonella typhimurium.
-Each compound is modeled as a graph, where nodes represent atoms with categorical labels (encoded as one-hot vectors) based on atom types, and edges depict the chemical bonds connecting them.
-The dataset comprises 188 molecular graphs with 7 unique node types.
-  • BZR [^7] https://chrsmrrs.github.io/datasets/docs/datasets/ (Rossi and Ahmed, 2015) consists of 405 molecular graphs representing ligands that interact with benzodiazepine receptors.
-Each molecule is treated as an independent graph and is classified into one of two categories.
-  • COX2888https://chrsmrrs.github.io/datasets/docs/datasets/ (Rossi and Ahmed, 2015) includes 467 molecular structures of cyclooxygenase-2 inhibitors.
-In this dataset, nodes correspond to atoms, while edges define chemical bonds—which may be single, double, triple, or aromatic.
-The molecules are divided into two distinct classes.
-  • PROTEINS [^9] https://chrsmrrs.github.io/datasets/docs/datasets/ (Borgwardt et al., 2005) is a dataset of protein structure graphs that encode both biochemical and structural properties.
-In this dataset, nodes represent secondary structural elements, while edges capture connectivity based on spatial proximity or amino acid sequence adjacency.
-Each node falls into one of three categories, and graphs are classified into two broader groups.
-B. Further Descriptions of Baselines
-In this section, we provide additional details about the baselines used in our experiments.
-(1) Supervised GNNs:
-  • GCN (Kipf and Welling, 2017): A graph neural network that aggregates node information using mean-pooling, thereby enabling nodes to capture structural information from their neighbors.
-  • GAT (Veličković et al., 2018): Unlike GCN, GAT incorporates attention mechanisms to assign different weights to neighboring nodes, refining the aggregation process based on their relative importance.
-(2) Graph Pre-training and Fine-tuning Models:
-  • DGI (Veličković et al., 2018): A self-supervised pre-training method that maximizes mutual information between local node embeddings and the graph’s global representation, thereby enhancing structural awareness.
-  • InfoGraph (Sun et al., 2020): An extension of DGI designed for graph-level tasks, aligning node and graph representations by optimizing their similarity.
-  • GraphCL (You et al., 2020): A contrastive learning framework that leverages diverse graph augmentations to extract structural patterns, aiming to improve representation consistency across transformations.
-(3) Single-step Prompting Models:
-  • ProG (Sun et al., 2023): Reformulates node- and edge-level tasks as graph-level problems by employing prompt graphs with task-specific structures to guide adaptation.
-  • GPF (Fang et al., 2024): A universal prompt-tuning strategy for pre-trained graph models that transforms input graph features to mimic various prompting effects.
-  • GPF+ (Fang et al., 2024): An enhanced version of GPF that integrates an attention mechanism to dynamically refine prompt representations.
-  • GraphPrompt (Liu et al., 2023): Bridges pre-training and downstream tasks using subgraph similarity-based prompting, where a learnable prompt is optimized to incorporate task-relevant information for both node and graph classification.
-C. Implementation Details
-Environment. Optimizer: The environment in which we ran all experiments is listed below.
-  • Operating system: Ubuntu 22.04.2,
-  • CPU information: AMD EPYC 7742 64-Core Processor,
-  • GPU information: NVIDIA GeForce RTX 3090 (24 GB).
+We adopt accuracy as the performance metric and report both the mean and standard deviation across these runs, in line with previous studies (Wang et al., 2020; Liu et al., 2021a, 2023).
+5.2. Implementation Details
 We outline key settings for the baselines and GCoT.
-Baseline settings. We utilized the official codes for all open-source baselines. Each model was tuned based on the settings recommended in their respective work to achieve optimal performance.
+Baseline settings. We utilize the official codes for all open-source baselines. Each model was tuned based on the settings recommended in their respective work to achieve optimal performance.
 For the baseline GCN (Kipf and Welling, 2017), we employ a 2-layer architecture, and set the hidden dimensions to 64.
@@
 For GraphPrompt (Liu et al., 2023), a 3-layer GCN is used as the backbone for all datasets, with the hidden dimensions set to 256.
-For GPF (Fang et al., 2024), we employ a 3-layer GCN as the backbone for all datasets, following the recommended settings.
+For GPF and GPF+ (Fang et al., 2024), we employ a 3-layer GCN as the backbone for all datasets.
 The hidden dimensions are set to 256.
-For GPF+ (Fang et al., 2024), we employ a 3-layer GCN as the backbone for all datasets, following the recommended settings.
-The hidden dimensions are set to 256.
-For ProG (Sun et al., 2023),
+For ProG (Sun et al., 2023), we employ a 2-layer GCN as the backbone for all datasets.
+The hidden dimensions are set to 100.
 GCoT settings.
 For our proposed GCoT, we utilize a 3-layer GCN as the backbone for all datasets, with the hidden dimensions set to 256.
-We set the number of inference step as 2, and the number of bias prompts to generate inital prompts as 5.
-D. Details about Heterophilic Datasets
-To examine the robustness of GCoT on heterophilic graphs, we conduct experiments on heterophilic datasets in Sect.
-5.4.
-The details of these datasets are as follows: (1) Cornell (Pei et al., 2020) is a network of 183 nodes representing webpages, with 295 edges indicating hyperlink connections among them.
+We set the number of inference step as 2 for node classification task, and 3 for graph classification task.
+The condition-net is implemented as a two-layer MLP with a bottleneck architecture (Yu et al., 2025c, b).
+Its input dimension is 256, while the hidden dimension is set to 32 for node classification and 8 for graph classification.
+Environment. All experiments were conducted on Ubuntu 22.04.2, using a machine equipped with AMD EPYC 7742 64-core processors and NVIDIA GeForce RTX 3090 (24 GB) GPUs.
+5.3. Performance Evaluation
+We first evaluate one-shot classification tasks.
+Then, we examine the model performance as the number of shots increases.
+Note that for the other experiments in Sect.
+5.4 and thereafter, we adopt the one-shot setting.
+One-shot performance. We present the results for one-shot node and graph classification in Table 2. We make several major observations, as follows.
+(1) GCoT consistently outperforms the baseline methods across both node and graph classification, demonstrating its overall advantage and robustness.
+(2) Supervised methods tend to underperform compared to others, as they do not leverage any pre-trained model.
+On the other hand, graph pre-training models achieve improved performance through self-supervised pre-training on unlabeled graphs.
+(3) Standard graph prompt learning approaches—ProG, GPF, GPF+, and GraphPrompt—often outperform fine-tuning of the pre-trained models, due to their alignment between pre-training and downstream objectives, as well as their parameter efficiency.
+However, they still fall short compared to GCoT, due to their single-step inference process that lacks iterative refinement of the final answer.
+This highlights the effectiveness of our CoT-style prompting, which enables step-by-step inference.
+Few-shot performance.
+To investigate the effect of labeled data size on performance, we vary the number of shots in both node and graph classification tasks.
+The results are presented in Fig.
+3, comparing GCoT against several competitive baselines.
+We make the following observations.
+(1) GCoT generally outperforms the baselines on both node and graph classification tasks, particularly in low-shot settings (e.g., $m\leq 5$), where labeled data are limited.
+(2) As the number of shots increases (e.g., $m>5$), all methods generally exhibit improved performance, as expected.
+Nevertheless, GCoT remains highly competitive, often achieving the best or near-best results.
+(3) On the PROTEINS dataset, the performance of all methods tends to fluctuate more.
+A possible reason is that this dataset exhibits greater variability in graph sizes compared to other datasets: The standard deviation of graph sizes in PROTEINS is 45.78, whereas other datasets fall within the range of 4.04 to 15.29.
+This may contribute to the unstable performance.
+Despite this, GCoT demonstrates greater robustness than the competing methods.
+Figure 3. Impact of labeled data size (number of shots) on node and graph classification.
+5.4. Ablation Study and Visualization
+To thoroughly evaluate the contribution of CoT-style prompting to graph models, we perform an ablation study and visualization.
+[…]
+[…]
+[…]
+Table 3. Ablation study on the effects of key components.
+Methods Node classification Graph classification
+Cora Pubmed MUTAG PROTEINS
+[…]
+GCoT-L1 57.18±14.34 63.31±10.05 56.54±14.12 54.71±8.57
+GCoT-L2 57.00±14.48 63.20±10.08 57.68±13.84 54.77±8.81
+GCoT-L3 57.01±14.66 63.33±10.05 57.85±16.10 56.22±8.45
+GCoT 59.67±15.51 64.43± 9.96 58.75±15.42 56.24±8.60
+Figure 4. Visualization of node and thought embeddings. Different colors represent different node classes.
+5.5. Heterophily Sensitivity
+To examine the robustness of GCoT on heterophilic graphs (Yu et al., 2025c), we conduct one-shot node classification on Wisconsin (Pei et al., 2020) and Squirrel (Rozemberczki et al., 2021).
+Specifically, Wisconsin is a network of 251 nodes representing webpages, with 199 edges indicating hyperlink connections among them.
 Node features are constructed using a bag-of-words approach based on webpage content.
-Each node is assigned to one of five categories: student, project, course, staff, or faculty.
-(2) Squirrel (Rozemberczki et al., 2021) consists of 5,201 Wikipedia pages discussing predefined topics.
-Nodes represent individual webpages, while edges capture 217,073 hyperlink connections between them, forming a page-page network.
-The dataset is categorized into five groups based on average monthly traffic.
+Squirrel (Rozemberczki et al., 2021) consists of 5,201 Wikipedia pages discussing predefined topics.
+Nodes represent individual webpages, while edges capture 217,073 hyperlink connections between them.
 Node features are constructed using key informative nouns extracted from the text content of the Wikipedia pages.
+As shown in Table 4, GCoT outperforms other competitive baselines on heterophilic node classification.
+These results indicate that our CoT-style prompting can generalize across both homophilic and heterophilic graphs.
+Table 4. Accuracy (%) evaluation of node classification on heterophilic graphs.
+Methods Wisconsin Chameleon
+DGI 28.04 $\pm$ 06.47 19.33 $\pm$ 4.57
+GraphCL 29.85 $\pm$ 08.46 27.16 $\pm$ 4.31
+ProG 28.99 $\pm$ 11.14 25.90 $\pm$ 3.95
+GPF 32.18 $\pm$ 11.36 25.98 $\pm$ 3.71
+GPF+ 32.02 $\pm$ 11.17 26.12 $\pm$ 3.75
+GraphPrompt 31.48 $\pm$ 05.18 25.36 $\pm$ 3.99
+GCoT 32.80 $\pm$ 11.59 27.96 $\pm$ 3.89
+5.6. Flexibility of Graph Prompting Methods
+To evaluate the flexibility and robustness of GCoT, we evaluate its performance using various standard graph prompting methods as the standard prompt.
+We integrate ProG (Sun et al., 2023), GPF (Fang et al., 2024), GPF+ (Fang et al., 2024), and GraphPrompt (Liu et al., 2023) into our framework.
+Specifically, ProG, GPF, and GPF+ are employed to modify the input features at the first inference step, whereas GraphPrompt is utilized to modify the output embeddings at the final step.
+The results for both node and graph classification on four datasets are presented in Table 5.
+Across all cases, GCoT outperforms its standard prompting counterparts, highlighting the robustness and flexibility of our CoT design.
+Table 5. Accuracy (%) evaluation of GCoT with different standard prompting methods.
+Prompting Node classification Graph classification
+Cora Pubmed MUTAG PROTEINS
+ProG 50.59±14.64 63.07±11.96 51.99±14.50 52.73±6.57
+with GCoT 52.62±15.27 64.01±11.27 58.42±15.39 55.82±8.79
+GPF 57.60±13.88 55.63±10.96 56.55±13.95 53.05±7.62
+with GCoT 58.98±15.28 62.63±10.01 57.66±13.11 56.98±8.30
+GPF+ 57.42±13.87 57.16±10.99 56.81±12.93 54.58±8.70
+with GCoT 59.35±15.37 62.18±10.49 58.54±13.29 55.45±9.20
+GraphPrompt 54.25± 9.38 63.11±10.01 55.44±12.56 53.80±7.93
+with GCoT 58.40±14.74 63.79± 9.40 58.28±15.47 55.32±8.43
+In each group, the first row represents a standard prompting method, and the second row represents the standard prompting coupled with our CoT design.
+5.7. Hyperparameter Analysis
+Next, we study the sensitivity of two key hyperparameters, namely, the hidden dimension of the condition-net, and the number of inference steps performed.
+Hidden dimension of condition-net.
+In our experiments, we implement the condition-net as a two-layer MLP with a bottleneck design.
+To examine the effect of the MLP architecture, we vary its hidden dimension $s$ and present the results in Fig.
+5.
+The results reveal that $s=32$ generally yields strong results for node classification, and $s=8$ for graph classification.
+A smaller $s$ may restrict the model’s representational capacity, limiting its effectiveness, whereas a larger $s$ introduces additional trainable parameters and increases the risk of overfitting in few-shot scenarios.
+Figure 5. Impact of hidden dimension $s$ in the condition-net.
+Number of inference steps.
+We further investigate the influence of the number of inference steps, $K$, with the results summarized in Fig.
+6.
+For node classification, we observe that $K=2$ generally yields the best performance, whereas $K=3$ is optimal for graph classification.
+This discrepancy may stem from the fact that graph classification relies more on global structural information, which may be better captured with more inference steps.
+Notably, for the Proteins dataset, $K=7$ achieves the highest performance, likely due to its intrinsic complexity and variability, as discussed in Sect.
+5.3.
+Figure 6. Impact of number of inference steps $K$.
+5.8. Computational Efficiency Analysis
+To evaluate the efficiency of GCoT, we perform one-shot node classification on a large-scale dataset called CLUSTER (Dwivedi et al., 2023), which comprises 1,172,320 nodes and 43,038,630 edges.
+The accuracy evaluation is shown in Table 6, whereas inference time comparisons with strong baselines are reported in Table 7.
+Table 6. Accuracy (%) evaluation of node classification on the large-scale CLUSTER dataset.
+Methods DGI GraphCL GraphPrompt GPF+ GCoT
+Accuracy 21.27 23.82 22.18 24.27 27.54
+While GCoT maintains higher accuracy on the large-scale dataset, we observe additional inference time due to its multi-step inference on the test set.
+Such a limitation of CoT has also been observed in NLP.
+In our case, inference time increases roughly linearly with the number of steps, which is acceptable considering the significant accuracy gains.
+Table 7. Inference time (milliseconds) on the test set of a node classification task.
+Methods Cora Pubmed MUTAG CLUSTER
+GraphPrompt 3.90 5.62 5.85 955.25
+GPF+ 5.25 5.92 4.79 713.08
+GCoT 7.95 13.95 11.44 2340.00
+6. Conclusions
+In this paper, we proposed GCoT, the first CoT prompting framework for text-free graphs.
+Specifically, we introduced the notion of step-by-step inference, where each inference step consists of three substages: prompt-based inference, thought construction, and thought-conditioned prompt learning.
+Concretely, we first feed the prompt-modified query graph into a pre-trained graph encoder, and then construct a thought by fusing the hidden embeddings from all layers of the pre-trained encoder to capture hierarchical structural knowledge.
+Subsequently, conditioned on the thought, we generate a series of node-specific prompts to guide the next inference step.
+After multiple such inference steps, GCoT makes a prediction on the final “answer.” Lastly, we conducted extensive experiments on eight public datasets, demonstrating that GCoT significantly outperforms a range of state-of-the-art baselines.
+Acknowledgments
+This research / project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Proposal ID: T2EP20122-0041).
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
+Appendices
+A. Further Descriptions of Datasets
+  • Cora (McCallum et al., 2000) is a citation network composed of 2,708 research papers in the field of computing, each classified into one of seven categories.
+The network consists of 5,429 citation links between papers.
+Each paper is represented by a binary word vector, where each entry indicates the presence or absence of a word from a predefined vocabulary of 1,433 unique terms.
+  • Citeseer (Sen et al., 2008) includes 3,312 computer science publications, categorized into six distinct classes, separate from those in Cora.
+The citation network contains 4,732 edges.
+Each document is encoded as a binary word vector that captures the presence or absence of words from a dictionary comprising 3,703 unique terms.
+  • PubMed (Sen et al., 2008) is a citation network of 19,717 biomedical articles related to diabetes, divided into three categories.
+The network includes 44,338 citation edges.
+Unlike Cora and Citeseer, each document is represented by a TF/IDF-weighted word vector derived from a dictionary of 500 unique terms.
+  • Photo (Shchur et al., 2018) consists of 7,487 photography-related products, each assigned to one of eight categories.
+The co-purchase network contains 119,043 edges, where connections indicate products frequently bought together.
+Each product is described by a feature vector extracted from metadata and customer reviews, with category labels corresponding to product types.
+  • MUTAG (Debnath et al., 1991) is a dataset of nitroaromatic compounds aimed at predicting their mutagenic effects on Salmonella typhimurium.
+Each compound is modeled as a graph, where nodes represent atoms with categorical labels (encoded as one-hot vectors) based on atom types, and edges depict the chemical bonds connecting them.
+The dataset comprises 188 molecular graphs with 7 unique node types.
+  • BZR (Rossi and Ahmed, 2015) consists of 405 molecular graphs representing ligands that interact with benzodiazepine receptors.
+Each molecule is treated as an independent graph and is classified into one of two categories.
+  • COX2 (Rossi and Ahmed, 2015) includes 467 molecular structures of cyclooxygenase-2 inhibitors.
+In this dataset, nodes correspond to atoms, while edges define chemical bonds—which may be single, double, triple, or aromatic.
+The molecules are divided into two distinct classes.
+  • PROTEINS (Borgwardt et al., 2005) is a dataset of protein structure graphs that encode both biochemical and structural properties.
+In this dataset, nodes represent secondary structural elements, while edges capture connectivity based on spatial proximity or amino acid sequence adjacency.
+Each node falls into one of three categories, and graphs are classified into two broader groups.
+B. Further Descriptions of Baselines
+We provide additional details about the baseline methods used in our experiments.
+(1) Supervised GNNs.
+  • GCN (Kipf and Welling, 2017): A graph neural network that aggregates node information using mean-pooling, thereby enabling nodes to capture structural information from their neighbors.
+  • GAT (Veličković et al., 2018): Unlike GCN, GAT incorporates attention mechanisms to assign different weights to neighboring nodes, refining the aggregation process based on their relative importance.
+(2) Graph Pre-training Models.
+  • DGI (Veličković et al., 2018): A self-supervised pre-training method that maximizes mutual information between local node embeddings and the graph’s global representation, thereby enhancing structural awareness.
+  • InfoGraph (Sun et al., 2020): An extension of DGI designed for graph-level tasks, aligning node and graph representations by optimizing their similarity.
+  • GraphCL (You et al., 2020): A contrastive learning framework that leverages diverse graph augmentations to extract structural patterns, aiming to improve representation consistency across transformations.
+(3) Standard Graph Prompting Models.
+  • ProG (Sun et al., 2023): Reformulates node- and edge-level tasks as graph-level problems by employing prompt graphs with task-specific structures to guide adaptation.
+  • GPF (Fang et al., 2024): A universal prompt-tuning strategy for pre-trained graph models that transforms input graph features to mimic various prompting effects.
+  • GPF+ (Fang et al., 2024): An enhanced version of GPF that integrates an attention mechanism to dynamically refine prompt representations.
+  • GraphPrompt (Liu et al., 2023): Bridges pre-training and downstream tasks using subgraph similarity-based prompting, where a learnable prompt is optimized to incorporate task-relevant information for both node and graph classification.